### v2

# Density estimation using the perceptron

Abstract
We propose a new density estimation algorithm. Given $n$ i.i.d. samples from a distribution belonging to a class of densities on $\mathbb{R}^{d}$, our estimator outputs any density in the class whose “perceptron discrepancy” with the empirical distribution is at most $O(\sqrt{d/n})$. The perceptron discrepancy between two distributions is defined as the largest difference in mass that they place on any halfspace of $\mathbb{R}^{d}$. It is shown that this estimator achieves expected total variation distance to the truth that is almost minimax optimal over the class of densities with bounded Sobolev norm and Gaussian mixtures. This suggests that regularity of the prior distribution could be an explanation for the efficiency of the ubiquitous step in machine learning that replaces optimization over large function spaces with simpler parametric classes (e.g. in the discriminators of GANs).
We generalize the above to show that replacing the “perceptron discrepancy” with the generalized energy distance of Székeley-Rizzo [SR13] further improves total variation loss. The generalized energy distance between empirical distributions is easily computable and differentiable, thus making it especially useful for fitting generative models. To the best of our knowledge, it is the first example of a distance with such properties for which there are minimax statistical guarantees.
1 Introduction
A standard step in many machine learning algorithms is to replace an (intractable) optimization over a general function space with an optimization over a large parametric class (most often neural networks). This is done in supervised learning for fitting classifiers, in variational inference [BKM17, ZBKM18] for applying ELBO, in variational autoencoders [KW19] for fitting the decoder, in Generative Adversarial Networks (GANs) [${}^{+}$14, ACB17] for fitting the discriminator, in diffusion models [${}^{+}$20, ${}^{+}$22] for fitting the score function, and many other settings.
To be specific, let us focus on the example of GANs, which brought about the new era of density estimation in high-dimensional spaces. The problem setting is the following. We are given access to an i.i.d. data $X_{1},\dots,X_{n}\in\mathbb{R}^{d}$ sampled from an unknown distribution $\nu$ and a class of distributions $\mathcal{G}$ on $\mathbb{R}^{d}$ (the class of available “generators”). The goal of the learner is to find $\operatorname*{arg\,min}_{\nu^{\prime}\in\mathcal{G}}D(\nu^{\prime},\nu)$, where $D$ is some dissimilarity measure (“metric”) between probability distributions. In the case of GANs this measure is the Jensen-Shannon divergence $\mathrm{JS}(p,q)\triangleq\mathrm{KL}(p\|{1\over 2}p+{1\over 2}q)+\mathrm{KL}(q\|{1\over 2}p+{1\over 2}q)$ where $KL(p\|q)=\int p(x)\log{p(x)\over q(x)}dx$ is the Kullback-Leibler divergence. As any $f$-divergence, JS has a variational form, cf. [PW23, Example 7.5]: $\mathrm{JS}(p,q)=\log 2+\sup_{h:\mathbb{R}^{d}\to(0,1)}\mathbb{E}_{p}[h]+\mathbb{E}_{q}[\log(1-h)]\,.$ With this idea in mind, we can now restate the objective of minimizing $\mathrm{JS}(\nu^{\prime},\nu)$ as a game between a “generator” $\nu^{\prime}$ and a “discriminator” $h$, i.e., the GAN’s estimator is
$\tilde{\nu}\in\operatorname*{arg\,min}_{\nu^{\prime}}\sup_{h:\mathbb{R}^{d}\to(0,1)}{1\over n}\sum_{i=1}^{n}h(X_{i})+\mathbb{E}_{\nu^{\prime}}[\log(1-h)]\,,$ (1.1)
where we also replaced the expectation over (the unknown) $\nu$ with its empirical version $\nu_{n}\triangleq\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$. Subsequently, the idea was extended to other types of metrics, notably the Wasserstein-GAN [ACB17], which defines
$\tilde{\nu}\in\operatorname*{arg\,min}_{\nu^{\prime}\in\mathcal{G}}\sup_{f\in\mathcal{D}}\left|\mathbb{E}_{Y\sim\nu^{\prime}}f(Y)-\frac{1}{n}\sum_{i=1}^{n}f(X_{i})\right|,$ (1.2)
where the set of discriminators $\mathcal{D}$ is a class of Lipschitz functions (corresponding to the variational characterization of the Wasserstein-1 distance).
The final step to turn (1.1) or (1.2) into an algorithm is to relax the domain of the inner maximization (“discriminator”) to a parametric class of neural network discriminators $\mathcal{D}$. Note that replacing $\sup_{h:\mathbb{R}^{d}\to(0,1)}$ with $\sup_{h\in\mathcal{D}}$ effectively changes the objective from minimizing the JS divergence to minimizing a “neural-JS”, similar to how MINE [${}^{+}$18] replaces the true mutual information with a “neural” one. This weakening is quite worrisome for a statistician. Indeed, while the JS divergence is a strong statistical distance (for example, it bounds total variation from above and from below [PW23, (7.39)]), the “neural-JS” is unlikely to possess any such properties.
How does one justify this restriction to a simpler class $\mathcal{D}$? A practitioner would say that while taking $\max_{h\in\mathcal{D}}$ restricts the power of the discriminator, the design of $\mathcal{D}$ is fine-tuned to picking up those features of the distributions that are relevant to the human eye [^1] Implying that whether or not total variation $\mathsf{TV}(\tilde{\nu},\nu)$ is high is irrelevant as long as the generated images look “good enough” to humans.. A theoretician, instead, would appeal to universal approximation results about neural networks to claim that restriction to $\mathcal{D}$ is almost lossless.
The purpose of this paper is to suggest (and prove) a third explanation: the answer is in the regularity of $\nu$ itself. Indeed, we show that the restriction of discriminators to a very small class $\mathcal{D}$ in (1.1) results in almost no loss of (minimax) statistical guarantees, even if $\mathcal{D}$ is far from being a universal approximator. That is, the minimizing distribution $\tilde{\nu}$ selected with respect to a weak form of the distance enjoys almost minimax optimal guarantees with respect to the strong total variation distance, provided that the true distribution $\nu$ is regular enough. Phrased yet another way, even though the “neural” distance is very coarse and imprecise, and hence the minimizer selected with respect to it might be expected to only fool very naive discriminators, in reality it turns out to fool all (arbitrarily complex, but bounded) discriminators.
Let us proceed to a more formal statement of our results. One may consult Section 1.3 for notation. We primarily focus on two classes of distributions on $\mathbb{R}^{d}$: first, $\mathcal{P}_{S}(\beta,d,C)$ denotes the set of distributions supported on the $d$-dimensional unit ball $\mathbb{B}(0,1)$ that have a density with finite $L^{2}$ norm and whose $(\beta,2)$-Sobolev norm (defined in (1.7)) is bounded by $C$; second, $\mathcal{P}_{G}(d)=\{\mu*\mathcal{N}(0,1):\operatorname{supp}(\mu)\subseteq\mathbb{B}(0,1)\}$ is the (non-parametric) class of Gaussian mixtures with compactly supported mixing distribution. We remind the reader that the total variation distance has the variational form $\mathsf{TV}(p,q)=\sup_{h:\mathbb{R}^{d}\to[0,1]}\mathbb{E}_{p}h-\mathbb{E}_{q}h\,.$ Our first result concerns the following class of discriminators:
$\mathcal{D}_{1}=\{x\mapsto\mathbbm{1}\{x^{\top}v\geq b\}:v\in\mathbb{R}^{d},b\in\mathbb{R}\}\,,$
the class of affine classifiers, which can be seen as a single layer perceptron with a threshold non-linearity.
Theorem 1.
For any $\beta>0$, $d\geq 1$ and $C>0$, there exists a finite constant $C_{1}$ so that
$\sup_{\nu\in\mathcal{P}_{S}(\beta,d,C)}\mathbb{E}\mathsf{TV}(\tilde{\nu},\nu)\leq C_{1}n^{-\frac{\beta}{2\beta+d+1}}\,,$ (1.3)
where the estimator $\tilde{\nu}$ is defined in (1.2) with $\mathcal{D}=\mathcal{D}_{1}$ and $\mathcal{G}=\mathcal{P}_{S}(\beta,d,C)$. Similarly, for any $d\geq 1$ there exists a finite constant $C_{2}$ so that
$\sup_{\nu\in\mathcal{P}_{G}(d)}\mathbb{E}\mathsf{TV}(\tilde{\nu},\nu)\leq C_{2}\frac{(\log(n))^{\frac{2d+2}{4}}}{\sqrt{n}}\,,$
where the estimator $\tilde{\nu}$ is defined in (1.2) with $\mathcal{D}=\mathcal{D}_{1}$ and $\mathcal{G}=\mathcal{P}_{G}(d)$.
Recall the classical result  [IK83] which shows that the minimax optimal estimation rate in TV over the class $\mathcal{P}_{S}(\beta,d,C)$ equals $n^{-\beta/(2\beta+d)}$ (up to constant factors). Thus, the estimator in (1.3) is almost optimal, the only difference being that the dimension $d$ is replaced by $d+1$. Similarly, for the Gaussian mixtures we reach the parametric rate up to a polylog factor [^2] For estimation of Gaussian mixtures in total variation the precise value of the minimax optimal polylog factor is at present unknown. However, for the $L_{2}$ distance the minimax rate is known, and in the course of our proofs (see (3.5)) we show that our estimator only loses a multiplicative factor of $\log(n)^{1/4}$ in loss compared to the optimal $L_{2}$-rate $\log(n)^{d/4}/\sqrt{n}$ derived in [KG22]..
The proof of Theorem 1 relies on a comparison inequality between total variation and the “perceptron discrepancy”, or maximum halfspace distance, which we define as
$\overline{d_{H}}(\mu,\nu)\triangleq\sup_{f\in\mathcal{D}_{1}}\{\mathbb{E}_{\mu}f-\mathbb{E}_{\nu}f\}.$
Note first that $\overline{d_{H}}\leq\mathsf{TV}$ clearly holds since all functions in the class $\mathcal{D}_{1}$ are bounded by $1$. For the other direction, by proving a generalization of the Gagliardo-Nirenberg-Sobolev inequality we derive the following comparisons.
Theorem 2.
For any $\beta>0$, $d\geq 1$ and $C>0$, there exists a finite constant $C_{1}$ so that
$\mathsf{TV}(\mu,\nu)^{\frac{2\beta+d+1}{2\beta}}\leq C_{1}\overline{d_{H}}(\mu,\nu)$ (1.4)
holds for all $\mu,\nu\in\mathcal{P}_{S}(\beta,d,C)$. Similarly, for any $d\geq 1$ there exists a finite constant $C_{2}$ such that
$\,\mathsf{TV}(\mu,\nu)\log\left(3+\frac{1}{\mathsf{TV}(\mu,\nu)}\right)^{-\frac{d+1}{2}}\leq C_{2}\overline{d_{H}}(\mu,\nu)$
holds for all $\mu,\nu\in\mathcal{P}_{G}(d)$.
We remark that we also show (in Proposition 7) that the exponent $2\beta+d+1\over 2\beta$ in (1.4) is tight, i.e. cannot be improved in general.
With Theorem 2 in hand the proof of Theorem 1 is notably simple. For example, let us prove (1.3) (for full details, see Section 4.2). Recall that $X_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\nu$, $\nu_{n}$ is the empirical distribution and $\tilde{\nu}=\operatorname*{arg\,min}_{\nu^{\prime}\in\mathcal{P}_{S}}\overline{d_{H}}(\nu^{\prime},\nu_{n})$. We then have from the triangle inequality and minimality of $\tilde{\nu}$:
$\overline{d_{H}}(\tilde{\nu},\nu)\leq\overline{d_{H}}(\tilde{\nu},\nu_{n})+\overline{d_{H}}(\nu_{n},\nu)\leq 2\overline{d_{H}}(\nu_{n},\nu)\,.$
Thus, from Theorem 2 we have
$\mathsf{TV}(\tilde{\nu},\nu)\leq(2C_{1}\overline{d_{H}}(\nu_{n},\nu))^{\frac{2\beta}{2\beta+d+1}}\,.$ (1.5)
Lastly, we recall that $\mathcal{D}_{1}$ is a class with finite VC-dimension and thus from uniform convergence (Theorem 8.3.23, [Ver18]) we have for some dimension-dependent constant $C$ that
$\mathbb{E}[\overline{d_{H}}(\nu_{n},\nu)]\leq{C\over\sqrt{n}}\,.$
Thus, applying expectation and Jensen’s inequality to (1.5) we get
$\mathbb{E}[\mathsf{TV}(\tilde{\nu},\nu)]\leq C\mathbb{E}[\overline{d_{H}}(\nu_{n},\nu)^{\frac{2\beta}{2\beta+d+1}}]\leq C\mathbb{E}[\overline{d_{H}}(\nu_{n},\nu)]^{\frac{2\beta}{2\beta+d+1}}\leq Cn^{-\frac{\beta}{2\beta+d+1}}$
as claimed.
While we believe that Theorem 1 provides theoretical proof for the efficacy of simple discriminators, it has several serious theoretical and practical deficiencies that we now address. First, the rate for the class $\mathcal{P}_{S}$ is not minimax optimal. In this regard, we show that by replacing the perceptron class $\mathcal{D}_{1}$ with a generalized perceptron
$\mathcal{D}_{\gamma}=\{x\mapsto|x^{\top}v-b|^{\gamma-1\over 2}:v\in\mathbb{R}^{d},b\in\mathbb{R}\}\,,\qquad\gamma\in(0,2)$
and taking an average over $\mathcal{D}_{\gamma}$ instead of a supremum, we are able to achieve a total variation rate of $n^{-\beta/(2\beta+d+\gamma)}$, thus coming arbitarily close to minimax optimality [^3] More precisely, within a polylog in $n$. as $\gamma\to 0$. See Theorem 6 for details.
Second, from the implementation point of view, the density estimation algorithm behind Theorem 1 is completely impractical. Indeed, finding the halfspace with maximal separation between even two empirical measures is a nonconvex, non-differentiable problem and takes super-poly time in the dimension $d$ assuming $\mathsf{P}\neq\mathsf{NP}$ [GR09], and $\omega(d^{\omega(\varepsilon^{-1})})$ time for $\varepsilon$-optimal agnostic learning between two densities assuming either SIVP or gapSVP [Tie23].
Even if we disregard the computational complexity, it is unclear how to find the exact minimizer $\tilde{\nu}$ of $\arg\min_{\nu^{\prime}}\overline{d_{H}}(\nu^{\prime},\nu_{n})$. This concern is alleviated by the fact that any $\tilde{\nu}$ satisfying $\overline{d_{H}}(\tilde{\nu},\nu_{n})=\mathcal{O}(\sqrt{d/n})$ will work withouth degrading our performance guarantee, and thus only an approximate minimizer is needed. Taking this one step further, our proof proceeds by replacing the perceptron discrepancy $\overline{d_{H}}$ (defined with respect to the best perceptron) with an average version $d_{H}$ defined in (2.2), for which the comparison in Theorem 2 still holds. Therefore, one does not even need to find an approximately optimal half-space, as random half-spaces provide sufficient discriminatory power.
Somewhat unexpectedly, we discover that the average perceptron discrepancy $d_{H}$ exactly equals Székely and Rizzo’s energy distance $\mathcal{E}_{1}$ (Definition 1, [SR13]), defined as
$\mathcal{E}_{1}^{2}(\mu,\nu)\triangleq\mathbb{E}\left[2\|X-Y\|-\|X-X^{\prime}\|-\|Y-Y^{\prime}\|\right],\qquad(X,X^{\prime},Y,Y^{\prime})\sim\mu^{\otimes 2}\otimes\nu^{\otimes 2}\,,$ (1.6)
[…]
From the algorithmic point of view our message is the following. If one has access to a parametric family of generators sampling from $\nu_{\theta}$ for parameters $\theta\in\mathbb{R}^{p}$, and if one can compute $\nabla_{\theta}$ of the generator forward pass, e.g., via pushforward of a reference distribution under a smooth transport map or neural network-based models [WM22, MRWZ23], then one can fit $\theta$ to the empirical sample $\nu_{n}$ by running stochastic gradient descent steps:
  • sample $m$ samples from $\nu_{\theta}$ and form the empirical distribution $\nu^{\prime}_{m}$,
  • compute the loss $\mathcal{E}_{1}(\nu^{\prime}_{m},\nu_{n})$ and backpropagate the gradient with respect to $\theta$,
  • update $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{E}_{1}(\nu^{\prime}_{m},\nu_{n})$ for some step size $\eta$.
Note that the computation of $\mathcal{E}_{1}(\nu^{\prime}_{m},\nu_{n})$ according to (1.6) requires $O(n^{2}+m^{2})$ steps and is friendly to gradient evaluations.
1.1 Contributions
To summarize, our main contributions are as follows. We show that $\beta$-smooth distributions, Gaussian mixtures and discrete distributions that are far apart in total variation distance must possess a halfspace on which their mass is substantially different (Theorems 2, 3 and 4).
We apply the separation results to density estimation problems, showing that an ERM density estimator nearly attains the minimax optimal density estimation rate with respect to $\mathsf{TV}$ over the aforementioned distribution classes (Theorems 1, 6 and 8).
In Section 2 we show that the average halfspace separation distance $d_{H}$ is equal up to constant to the energy distance $\mathcal{E}_{1}$ (Proposition 2), which has many equivalent expressions: as a weighted $L^{2}$-distance between characteristic functions (Proposition 3), as the sliced Cramér-$2$ distance (Proposition 4), as an IPM/MMD/energy distance (Section 2.3), and as the $L^{2}$-norm of the Riesz potential (Proposition 5).
We generalize the average halfspace distance $d_{H}$ to include an exponent $\gamma\in(0,2)$, corresponding to the generalized energy distance $\mathcal{E}_{\gamma}$. Consequently, we discover that if instead of thresholded linear features $\mathbbm{1}\{v^{\top}x>b\}$ we use the non-linearity $|v^{\top}x-b|^{\gamma}$, smooth distributions and Gaussian mixtures can be separated even better (Theorem 3). Combined with the fact that $\mathcal{E}_{\gamma}$, similarly to $d_{H}$, decays between population and sample measures at the parametric rate (Lemma 1), the ERM for $\mathcal{E}_{\gamma}$ reduces the slack in the density estimation rate, almost achieving minimax optimality. This result, combined with its strong approximation properties, supports its use in modern generative models (e.g. [HJA20, ${}^{+}$14, ${}^{+}$22, ${}^{+}$]).
Finally, Proposition 10 shows that recent work applying $\overline{d_{H}}$ for two-sample testing is sub-optimal over the class of smooth distributions in the minimax sense.
1.2 Related work
In the statistics literature, an estimator of the form (1.2) appears in the famous work [Yat85]. Instead of indicators of halfspaces, they consider the class of discriminators
$\mathcal{Y}_{\epsilon}\triangleq\{\mathbbm{1}\{\mathrm{d}\nu_{i}/\mathrm{d}\nu_{j}\geq 1\}:1\leq i,j\leq N(\epsilon,\mathcal{G})\},$
where $\nu_{1},\dots,\nu_{N(\epsilon,\mathcal{G})}$ forms a minimal $\epsilon$-$\mathsf{TV}$ covering of the class $\mathcal{G}$ and $N(\epsilon,\mathcal{G})$ is the so-called covering number. Writing $d_{Y}(\mu,\mu^{\prime})=\sup_{f\in\mathcal{Y}}(\mathbb{E}_{\mu}f-\mathbb{E}_{\mu^{\prime}}f)$, it is not hard to prove that $|\mathsf{TV}-d_{Y}|=O(\epsilon)$ on $\mathcal{G}\times\mathcal{G}$ and that $\mathbb{E}d_{Y}(\nu,\nu_{n})\lesssim\sqrt{\log N(\epsilon_{n},\mathcal{G})/n}$ by a union bound coupled with a binomial tail inequality. From here $\mathbb{E}\mathsf{TV}(\tilde{\nu},\nu)\lesssim\sqrt{\log N(\epsilon,\mathcal{G})/n}+\epsilon$ follows by the triangle inequality (here $\tilde{\nu}$ is defined as in (1.2) with $\mathcal{D}=\mathcal{Y}$). Note that in contrast to our perceptron discrepancy $\overline{d_{H}}$, Yatracos’ estimator attains the optimal rate on $\mathcal{G}=\mathcal{P}_{S}$, corresponding to the choice $\epsilon=\epsilon(n)\asymp n^{-\beta/(2\beta+d)}$.
Several other related works such as [${}^{+}$18, Lia21] study the problem of minimax density estimation over classical smoothness classes with respect to Integral Probability Metrics (IPMs) $d_{\mathcal{D}}(\mathbb{P},\mathbb{Q})\triangleq\sup_{f\in\mathcal{D}}|\mathbb{E}_{\mathbb{P}}f-\mathbb{E}_{\mathbb{Q}}f|$. In particular, these works seek estimators $\tilde{\nu}$ such that $d_{\mathcal{D}}(\tilde{\nu},\nu)$ is small for some discriminator $\mathcal{D}$. Note some crucial differences to our work: first, we evaluate performance with respect to total variation in Theorem 1 which bears more interest both theoretically and empirically; second, we restrict our attention to estimators $\tilde{\nu}$ attained by ERM which is more commonly used in practice, while corresponding results of [${}^{+}$18, Lia21] consider classical orthogonal projection estimators.
A paper closer in spirit to ours is [BMR18] whose authors study comparison inequalities between the Wasserstein distance $W_{1}$ and the IPM $d_{relu}$ defined by the discriminator class $\mathcal{D}=\{x\mapsto\operatorname{Relu}(x^{\top}v+b):b,\|v\|\leq 1\}$. They show [BMR18, Theorem 3.1] that $\sqrt{\kappa/d}W_{1}\lesssim d_{relu}\lesssim W_{1}$ for Gaussian distributions with mean in the unit ball, where $\kappa$ is an upper bound on their condition numbers and $d$ is the dimension. They obtain results for other distribution classes (Gaussian mixtures, exponential families), but for each of these they use a different class of discriminators that is adapted to the problem. In contrast, we mainly study the discriminator class $\mathcal{D}_{1}=\{x\mapsto\mathbbm{1}\{x^{\top}v\geq b\}:\|v\|\leq 1,b\in\mathbb{R}\}$ and are able to derive novel comparisons to $\mathsf{TV}$ for smooth distributions, Gaussian mixtures and discrete distributions. In addition, we prove the (near)-optimality of our results (for smooth densities) and also derive nonparametric estimation rates for the corresponding GAN density estimators.
Independent of this work, recent results by [PCGT23] investigate the halfspace separability of distributions for the setting of two-sample testing. However, their focus was on the asymptotic power of the test as the number of samples grows to infinity. Our lower bound construction presented in Appendix D proves that their proposed test is sub-optimal in the minimax setting. See Section 5 for a more detailed discussion.
1.3 Notation
The symbols $O,o,\Theta,\Omega,\omega$ follow the conventional “big-O” notation, and $\tilde{O},\tilde{o}$ hide polylogarithmic factors. We use $\lesssim$, $\gtrsim$ and $\asymp$ throughout our calculations to hide multiplicative constants that are irrelevant (depending on the context). Given a vector $x\in\mathbb{R}^{d}$, we write $\|x\|$ for its Euclidean norm and $\langle x,y\rangle\triangleq x^{\top}y$ for the Euclidean inner product of $x,y\in\mathbb{R}^{d}$. The Gamma function is denoted by $\Gamma$. We write $\mathbb{B}(x,r)\triangleq\{y\in\mathbb{R}^{d}:\|x-y\|\leq r\}$, $\mathbb{S}^{d-1}\triangleq\{x\in\mathbb{R}^{d}:\|x\|=1\}$ and $\sigma$ for the unnormalized surface measure on $\mathbb{S}^{d-1}$. The surface area of a unit $(d-1)$-sphere is also written as $\sigma(\mathbb{S}^{d-1})=2\pi^{d/2}/\Gamma{\bigl{(}}{\frac{d}{2}}{\bigr{)}}$. In particular, if $X$ is a random vector uniformly distributioned on $\mathbb{S}^{d-1}$ then for any $h$ we have
$\mathbb{E}[h(X)]=\frac{1}{\sigma(\mathbb{S}^{d-1})}\int_{\mathbb{R}^{d}}h(y)\mathrm{d}\sigma(y)\,.$
The convolution between functions/measures is denoted by ***. We write $L^{p}(\mathbb{R}^{d})$ for the space of (equivalence classes of) functions $\mathbb{R}^{d}\to\mathbb{C}$ that satisfy $\|f\|_{p}\triangleq\left(\int_{\mathbb{R}^{d}}|f(x)|^{p}\mathrm{d}x\right)^{1/p}<\infty$. The space of all probability distributions on $\mathbb{R}^{d}$ is denoted as $\mathcal{P}(\mathbb{R}^{d})$. For a signed measure $\nu$ we write $\operatorname{supp}(\nu)$ for its support and $M_{r}(\nu)\triangleq\int\|x\|^{r}\mathrm{d}|\nu|(x)$ for its $r$’th absolute moment. Given $\mathbb{P},\mathbb{Q}\in\mathcal{P}(\mathbb{R}^{d})$ we write $\mathsf{TV}(\mathbb{P},\mathbb{Q})\triangleq\sup_{A\subseteq\mathbb{R}^{d}}[\mathbb{P}(A)-\mathbb{Q}(A)]$ for the total variation distance, where the supremum is over all measurable sets.
Given a function $f\in L^{1}(\mathbb{R}^{d})$, define its Fourier transform as
$\widehat{f}(\omega)\triangleq\mathcal{F}[f](\omega)\triangleq\int_{\mathbb{R}^{d}}e^{-i\langle x,\omega\rangle}f(x)\mathrm{d}x.$
Given a finite signed measure $\nu$ on $\mathbb{R}^{d}$, define its Fourier transform as $\mathcal{F}[\nu](\omega)\triangleq\int_{\mathbb{R}^{d}}e^{-i\langle\omega,x\rangle}\mathrm{d}\nu(x)$. We extend the Fourier transform to $L^{2}(\mathbb{R}^{d})$ and tempered distributions in the standard manner. Given $f\in L^{2}(\mathbb{R}^{d})$ and $\beta>0$, define its homogenous Sobolev seminorm of order $(\beta,2)$ as
$\|f\|_{\beta,2}^{2}\triangleq\int_{\mathbb{R}^{d}}\|\omega\|^{2\beta}|\widehat{f}(\omega)|^{2}\mathrm{d}\omega.$ (1.7)
Further, we define two specific classes of functions of interest as follows: $\mathcal{P}_{S}(\beta,d,C)$ is a set of smooth densities while $\mathcal{P}_{G}(d)$ is a set of all Gaussian mixtures with support in the unit ball, formally
$\mathcal{P}_{S}(\beta,d,C)$$\triangleq\{\mu\in\mathcal{P}(\mathbb{R}^{d}):\operatorname{supp}(\mu)\subseteq\mathbb{B}(0,1),\text{$\mu$ has density $p$ with }\|p\|_{\beta,2}\leq C\},$
$\mathcal{P}_{G}(d)$$\triangleq\{\nu*\mathcal{N}(0,I_{d}):\nu\in\mathcal{P}(\mathbb{R}^{d}),\operatorname{supp}(\nu)\subseteq\mathbb{B}(0,1)\}.$
Assumption 1.
Throughout the paper we assume that $C$ in the definition of $\mathcal{P}_{S}(\beta,d,C)$ is large enough relative to $\beta$ and $d$, such that $\mathcal{P}_{S}(\beta,d,C/2)$ is non-empty.
1.4 Structure
The structure of the paper is as follows. In Section 2 we introduce the generalized energy distance, the main object of our study. We show how it relates to the perceptron discrepancy $\overline{d_{H}}$ and its relaxation $d_{H}$; we record equivalent formulations of the generalized energy distance, one of which is a novel “sliced-distance” form. In Section 3, we present our main technical results on comparison inequalities between total variation and the energy distance. In Section 4 we analyse the density estimator that minimizes the empirical energy distance, and prove Theorem 1 and Theorem 2 in Section 4.2. In Section 5 we show that the use of $\overline{d_{H}}$ for two sample testing results in suboptimal performance. We conclude in Section 6. All omitted proofs and auxiliary results are deferred to the Appendix.
2 The generalized energy distance
Given two probability distributions $\mu,\nu$ on $\mathbb{R}^{d}$ with finite $\gamma$’th moment, the generalized energy distance of order $\gamma\in(0,2)$ between them is defined as
$\mathcal{E}_{\gamma}(\mu,\nu)=\mathbb{E}\Big{[}2\|X-Y\|^{\gamma}-\|X-X^{\prime}\|^{\gamma}-\|Y-Y^{\prime}\|^{\gamma}\Big{]},\qquad\text{where }(X,X^{\prime},Y,Y^{\prime})\sim\mu^{\otimes 2}\otimes\nu^{\otimes 2}.$ (2.1)
As we alluded to in the introduction, the proof of Theorems 1 and 2 becomes possible once we relax the supremum in the definition of $\overline{d_{H}}$ to an unnormalized average over halfspaces. In Section 2.1 we discuss this relaxation in more detail and identify a connection to the energy distance $\mathcal{E}_{1}$ defined above in (2.1). Motivated by this, we study the (generalized) energy distance and give multiple equivalent characterizations of it from Section 2.1 to Section 2.5.
2.1 From perceptron discrepancy to energy distance
Our first goal is to connect the study of $\overline{d_{H}}$ to the study of $\mathcal{E}_{\gamma}$ with $\gamma=1$. To achieve this, we introduce an intermediary, the “average” perceptron discrepancy $d_{H}$. Given two probability distributions $\mu,\nu$ on $\mathbb{R}^{d}$, we define
$d_{H}(\mu,\nu)\triangleq\sqrt{\int_{v\in\mathbb{S}^{d-1}}\int_{b\in\mathbb{R}}\left(\int_{\langle v,x\rangle\geq b}\mathrm{d}\mu(x)-\mathrm{d}\nu(x)\right)^{2}\mathrm{d}{b}\mathrm{d}\sigma(v)},$ (2.2)
where $\sigma$ denotes the surface area measure.
If the two distributions $\mu,\nu$ are supported on a compact set, then the overall definition can indeed be regarded as a ‘mean squared’ version of perceptron discrepancy, because the integrals over $b$ and $v$ only range over bounded sets. However, in general, the integral over $b$ in the definition of $d_{H}$ is not normalizable and that is why we put “average” in quotes. Nevertheless, we have the following comparisons between $d_{H}$ and $\overline{d_{H}}$.
Proposition 1.
For any $\beta>0$, $d\geq 1$, $C>0$, and for all $\mu,\nu\in\mathcal{P}_{S}(\beta,d,C)$, we have
$\sqrt{\frac{\Gamma(d/2)}{4\pi^{d/2}}}d_{H}(\mu,\nu)\leq\overline{d_{H}}(\mu,\nu).$ (2.3)
Moreover, for all $d\geq 1$, there exists a finite constant $C_{1}$ such that for all $\mu,\nu\in\mathcal{P}_{G}(d)$,
$\frac{d_{H}(\mu,\nu)}{\log(3+1/d_{H}(\mu,\nu))^{1/4}}\leq C_{1}\overline{d_{H}}(\mu,\nu).$
Proof.
The proof of (2.3) is immediate after noting that all distributions in $\mathcal{P}_{S}(\beta,d,C)$ are supported on the $d$-dimensional unit ball and that $\int_{v\in\mathbb{S}^{d-1}}\int_{-1}^{1}\mathrm{d}b\mathrm{d}\sigma(v)=4\pi^{d/2}/\Gamma(d/2)$. Thus, we focus on the Gaussian mixture case. Write $\mu-\nu=\tau*\phi$ where $\phi$ denotes the density of the standard Gaussian $\mathcal{N}(0,I_{d})$ and $\tau\in\mathcal{P}(\mathbb{R}^{d})$ is the difference of the two implicit mixing measures. For any $R>0$, we have
$\overline{d_{H}}(\mu,\nu)$ $\geq\sup_{v\in\mathbb{S}^{d-1},|b|\leq R}\int_{\langle x,v\rangle\geq b}(\tau*\phi)(x)\mathrm{d}x$
$\geq\sqrt{\frac{1}{2R\operatorname{vol}_{d-1}(\mathbb{S}^{d-1})}\int_{\mathbb{S}^{d-1}}\int_{|b|\leq R}\left(\int_{\langle x,v\rangle\geq b}(\tau*\phi)(x)\mathrm{d}x\right)^{2}\mathrm{d}b\mathrm{d}\sigma(v)}.$
Now, since $\tau$ is supported on a subset of $\mathbb{B}(0,1)$ by definition of the class $\mathcal{P}_{G}(d)$, for any $v\in\mathbb{S}^{d-1}$ and $R\geq 2$ we have the bound
$\int_{|b|>R}\left(\int_{\langle x,v\rangle\geq b}\int_{\mathbb{R}^{d}}\phi(x-y)\mathrm{d}\tau(y)\mathrm{d}x\right)^{2}\mathrm{d}b$ $\leq\int_{|b|>R}\left(\int_{\langle x,v\rangle\geq|b|}\exp(-(\|x\|-1)^{2}/2)\mathrm{d}x\right)^{2}\mathrm{d}b$
$\leq\int_{|b|>R}\left(\int_{\|x\|\geq|b|}\exp(-\|x\|^{2}/8)\mathrm{d}x\right)^{2}\mathrm{d}b$
$\lesssim\exp(-\Omega(R^{2})),$
where we implicitly used that $\int\mathrm{d}\tau=0$ as $\tau$ is the difference of two probability distributions. Choosing $R\asymp\sqrt{\log(3+1/d_{H}(\mu,\nu))}$ concludes the proof. ∎
Proposition 1 implies that to obtain a comparison between $\mathsf{TV}$ and $\overline{d_{H}}$, specifically for lower bounding $\overline{d_{H}}$, it suffices to consider the relaxation $d_{H}$ instead. The next observation we make is that $d_{H}$ is in fact equal, up to constant, to the energy distance.
Proposition 2.
Let $\mu,\nu$ be probability distributions on $\mathbb{R}^{d}$ with finite mean. Then
$d_{H}(\mu,\nu)=\frac{\pi^{(d-1)/4}}{\sqrt{\Gamma(\frac{d+1}{2}})}\mathcal{E}_{1}(\mu,\nu).$
Proof.
This is a direct implication of (2.4) and (2.7). We defer the proof to the more general Proposition 4. ∎
As will be clear from the rest of the paper, it does pay off to study $\mathcal{E}_{\gamma}$ for general $\gamma$, even though so far we only justified its relevance to the results stated in the introduction for the case of $\gamma=1$. With this in mind, we proceed to study various properties of the generalized energy distances $\{\mathcal{E}_{\gamma}\}_{\gamma\in(0,2)}$.
2.2 The Fourier form
The formulation of the generalized energy distance that we rely on most heavily in our proofs is the following.
Proposition 3 ([SR13, Proposition 2]).
Let $\gamma\in(0,2)$ and let $\mu,\nu$ be probability distributions on $\mathbb{R}^{d}$ with finite $\gamma$’th moment. Then,
$\mathcal{E}_{\gamma}^{2}(\mu,\nu)=F_{\gamma}(d)\int_{\mathbb{R}^{d}}\frac{|\widehat{\mu}(\omega)-\widehat{\nu}(\omega)|^{2}}{\|\omega\|^{d+\gamma}}\mathrm{d}\omega,$ (2.4)
where we define $F_{\gamma}(d)=\frac{\gamma 2^{\gamma-1}\Gamma\left(\frac{d+\gamma}{2}\right)}{\pi^{d/2}\Gamma(1-\frac{\gamma}{2})}$.
Remark 1.
Note that $F_{\gamma}(d)=\Theta\left(\gamma(2-\gamma)\Gamma\left(\frac{d+\gamma}{2}\right)\pi^{-d/2}\right)$ up to a universal constant.
This shows that the generalized energy distance is a weighted $L^{2}$ distance in Fourier space. The fact that $\mathcal{E}_{\gamma}$ is a valid metric on probability distributions with finite $\gamma$’th moment is a simple consequence of Proposition 3. Another straightforward consequence of the above characterization is the fact that $\mathcal{E}_{\gamma}$ decays at the parametric rate between empirical and population measures. Recall that $M_{t}(\nu)$ denotes the $t$’th absolute moment of the measure $\nu$.
Lemma 1.
Let $\nu$ be a probability distribution on $\mathbb{R}^{d}$ and let $\nu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$ for an i.i.d. sample $X_{1},\dots,X_{n}$ from $\nu$. Then, for any $\gamma\in(0,2)$,
$\mathbb{E}\mathcal{E}_{\gamma}^{2}(\nu,\nu_{n})\leq\frac{10d^{\gamma/2}M_{\gamma}(\nu)}{n}.$
For a high-probability bound when $\nu$ is compactly supported, see Lemma 3.
Proof.
Let $X\sim\nu$. We use Proposition 3 as well as Tonelli’s theorem to interchange integrals to obtain
$\frac{1}{F_{\gamma}(d)}\mathbb{E}\mathcal{E}_{\gamma}^{2}(\nu,\nu_{n})$ $=\int_{\mathbb{R}^{d}}\frac{\mathbb{E}|\widehat{\nu}-\widehat{\nu}_{n}|^{2}}{\|\omega\|^{d+\gamma}}\mathrm{d}\omega$
$=\frac{1}{n}\int_{\mathbb{R}^{d}}\frac{\operatorname{var}(\cos\langle X,\omega\rangle)+\operatorname{var}(\sin\langle X,\omega\rangle)}{\|\omega\|^{d+\gamma}}\mathrm{d}\omega$
$\leq\frac{1}{n}\mathbb{E}\int_{\mathbb{R}^{d}}\frac{(\cos\langle X,\omega\rangle-1)^{2}+(\sin\langle X,\omega\rangle)^{2}}{\|\omega\|^{d+\gamma}}\mathrm{d}\omega$
[…]
Multiplying both sides by $F_{\gamma}(d)$ and using that $(2-\gamma)\Gamma(1-\gamma/2)\geq 1.77$ by the third point of Lemma 9 gives
$\mathbb{E}\mathcal{E}^{2}_{\gamma}(\nu,\nu_{n})\leq\frac{4.52\times 2^{\gamma}M_{\gamma}(\nu)}{n}\frac{\Gamma\left(\frac{d+\gamma}{2}\right)}{\Gamma\left(\frac{d}{2}\right)}.$
The conclusion now follows by the first point of Lemma 9. ∎
2.3 The MMD and IPM forms
Another interpretation of the generalized energy distance is through the theory of Maximum Mean Discrepancy (MMD). Given a set $\mathcal{X}$ and a positive semidefinite kernel $k:\mathcal{X}^{2}\to\mathbb{R}$, there is a unique reproducing kernel Hilbert space (RKHS) $\mathcal{H}_{k}$ consisting of the closure of the linear span of $\{k(x,\cdot),x\in\mathcal{X}\}$ with respect to the inner product $\left\langle k(x,\cdot),k(y,\cdot)\right\rangle_{\mathcal{H}_{k}}=k(x,y).$
For a probability distribution $\mu$ on $\mathcal{X}$, define its kernel embedding as $\theta_{\mu}=\int_{\mathbb{R}^{d}}k(x,\cdot)\mathrm{d}\mu(x)$. As shown in [${}^{+}$17, Lemma 3.1], the kernel embedding $\theta_{\mu}$ exists and belongs to the RKHS $\mathcal{H}_{k}$ if $\mathbb{E}[\sqrt{k(X,X^{\prime})}]<\infty$ for $(X,X^{\prime})\sim\mu^{\otimes 2}$ — as is the case for our kernel defined later in 2.5. Then, given two probability distributions $\mu$ and $\nu$, the MMD measures their distance in the RKHS by
$\operatorname{MMD}_{k}(\mu,\nu)\triangleq\|\theta_{\mu}-\theta_{\nu}\|_{\mathcal{H}_{k}}.$
We refer the reader to [SS01, ${}^{+}$17] for more details on the underlying theory. MMD has a closed form thanks to the reproducing property:
$\operatorname{MMD}_{k}^{2}(P,Q)$ $=\mathbb{E}\Big{[}k(X,X^{\prime})+k(Y,Y^{\prime})-2k(X,Y)\Big{]},$
where $(X,X^{\prime},Y,Y^{\prime})\sim\mu^{2}\otimes\nu^{2}$. Moreover, it also follows that MMD is an Integral Probability Metric (IPM) where the supremum is over the unit ball of the RKHS $\mathcal{H}_{K}$:
$\operatorname{MMD}_{k}(\mu,\nu)=\sup_{f\in\mathcal{H}_{k}:\|f\|_{\mathcal{H}_{k}}\leq 1}\mathbb{E}[f(X)-f(Y)].$
In our case, we can define the kernel
$k_{\gamma}(x,y)=\|x\|^{\gamma}+\|y\|^{\gamma}-\|x-y\|^{\gamma}$ (2.5)
for $\gamma\in(0,2)$, which in one dimension corresponds to the covariance operator of fractional Brownian motion. For a proof of the nontrivial fact that $k_{\gamma}$ above is positive definite see for example [SSGF13]. With the choice of $k_{\gamma}$ it follows trivially from its definition that the generalized energy distance $\mathcal{E}_{\gamma}$ is equal to the MMD with kernel $k_{\gamma}$, i.e.
$\mathcal{E}_{\gamma}(\mu,\nu)=\operatorname{MMD}_{k_{\gamma}}(\mu,\nu)$
for all distributions $\mu,\nu$ with finite $\gamma$’th moment. It is noteworthy that while $\overline{d_{H}}$ is by definition an IPM, so is its averaged version $d_{H}$.
2.4 The sliced form
Another equivalent characterization of the generalized energy distance is in the form of a sliced distance. Sliced distances are calculated by first choosing a random direction on the unit sphere, and then computing a one-dimensional distance in the chosen direction between the projections of the two input distributions. For $\gamma\in(0,2)$ define the function
$\psi_{\gamma}(x)=\begin{cases}|x|^{(\gamma-1)/2}&\text{for }\gamma\neq 1\\ \mathbbm{1}\{x\geq 0\}&\text{otherwise.}\end{cases}$ (2.6)
The following result, to the best of our knowledge, has not appeared in prior literature except for the case of $\gamma=1$.
Proposition 4.
Let $\gamma\in(0,2)$ and let $\mu,\nu$ be probability distributions on $\mathbb{R}^{d}$ with finite $\gamma$’th moment. Then for $(X,Y)\sim\mu\otimes\nu$ we have
$\mathcal{E}^{2}_{\gamma}(\mu,\nu)=\frac{1}{S_{\gamma}}\int_{\mathbb{S}^{d-1}}\int_{\mathbb{R}}\Big{[}\mathbb{E}\psi_{\gamma}(\langle X,v\rangle-b)-\mathbb{E}\psi_{\gamma}(\langle Y,v\rangle-b)\Big{]}^{2}\mathrm{d}b\mathrm{d}\sigma(v).$ (2.7)
where $S_{\gamma}=\frac{\pi^{\frac{d}{2}+1}\Gamma(1-\frac{\gamma}{2})}{\gamma 2^{\gamma-1}\Gamma(\frac{d+\gamma}{2})\cos^{2}(\frac{\pi(\gamma-1)}{4})\Gamma(\frac{1-\gamma}{2})^{2}}$ when $\gamma\neq 1$ and $S_{1}=\frac{\pi^{\frac{d-1}{2}}}{\Gamma(\frac{d+1}{2})}$.
The proof of Proposition 4 hinges on computing the Fourier transform of the function $\psi_{\gamma}$, which can be interpreted as a tempered distribution. We point out a special property of the integral on the right hand side of (2.7). After expanding the square, one finds that the individual terms in the sum are not absolutely integrable for $\gamma\neq 1$. However, due to cancellations within the squared quantity, the integral is finite.
As claimed, Proposition 4 allows us to interpret $\mathcal{E}_{\gamma}$ as a sliced probability divergence (using the language of [${}^{+}$20]). Given $v\in\mathbb{S}^{d-1}$, write $\theta_{v}=\langle v,\cdot\rangle$ and $\theta_{v}\#\nu=\nu\circ\theta_{v}$ for the pushforward of $\nu$ under $\theta_{v}$. We have
$S_{\gamma}(d)\mathcal{E}^{2}_{\gamma}(\mu,\nu)=S_{\gamma}(1)\int_{\mathbb{S}^{d-1}}\mathcal{E}^{2}_{\gamma}(\theta_{v}\#\mu,\theta_{v}\#\nu)\mathrm{d}\sigma(v).$
We may also observe that the energy distance $\mathcal{E}_{1}$ is equal to the sliced Cramér-$2$ distance [^4] The Cramér-$p$ distance is simply the $L^{p}$ distance between cumulative distribution functions. up to constant, which has been studied recently by both theoretical and empirical works [${}^{+}$18, KNSS22].
2.5 The Riesz potential form
The generalized energy distance can also be linked to the Riesz potential [Lan72, Chapter 1.1], which is the inverse of the fractional Laplace operator. Given $0<s<d$, the Riesz potential $I_{s}f$ of a compactly supported signed measure $f$ on $\mathbb{R}^{d}$ is defined (in a weak sense) by
${\displaystyle I_{s}f=f*K_{s}},$
where $K_{s}(x)=c_{s}^{-1}\|x\|^{s-d}$ and $c_{s}=\pi^{{d/2}}2^{s}\Gamma(s/2)\Gamma((d-s)/2)^{-1}$. The Fourier transform of the Riesz kernel is given by $\widehat{K_{s}}(\omega)=\|\omega\|^{{-s}}$, interpreted as a tempered distribution. The following proposition is derived by setting $s=\frac{d+\gamma}{2}$ and using the Fourier form (Proposition 3) of the energy distance.
Proposition 5.
Let $\gamma\in(0,\min\{d,2\})$ and let $\mu,\nu$ be compactly supported probability distributions on $\mathbb{R}^{d}$. Then
$\mathcal{E}_{\gamma}(\mu,\nu)=(2\pi)^{d/2}\sqrt{F_{\gamma}(d)}\|I_{\frac{d+\gamma}{2}}(\mu-\nu)\|_{2}.$ (2.8)
3 Main comparison: TV versus Energy
After considering the connection of the perceptron discrepancy $\overline{d_{H}}$ to the energy distance in Section 2, we turn to some of our main technical results, which provide novel quantitative comparisons between $\{\mathcal{E}_{\gamma}\}_{\gamma\in(0,2)}$ and the total variation distance. In Section 3.1 we show that the generalized energy distance is upper bounded by total variation for compactly supported distributions. In Section 3.2 we derive lower bounds on the generalized energy distance in terms of the total variation distance over the two distribution classes that we have introduced, namely smooth distributions and Gaussian mixtures. Finally, in Section 3.3 we turn to the case of discrete distributions, which requires alternative techniques.
3.1 Upper bound – compactly supported distributions
Note that we (obviously) always have $\overline{d_{H}}(\mu,\nu)\leq\mathsf{TV}(\mu,\nu)$ for arbitrary probability measures $\mu$ and $\nu$. Moreover, for distributions supported on a unit ball we also have $d_{H}(\mu,\nu)\lesssim\overline{d_{H}}(\mu,\nu)$. Therefore, by the identification of $d_{H}$ and $\mathcal{E}_{1}$ (Proposition 2), we can see that for distributions with bounded support, we always have $\mathcal{E}_{1}(\mu,\nu)\lesssim\mathsf{TV}(\mu,\nu)\,.$ The next result generalizes this estimate for all $\mathcal{E}_{\gamma}$, not just $\gamma=1$.
Proposition 6.
For any dimension $d\geq 1$ and $\gamma\in(0,\min\{d,2\})$ there exists a finite constant $c$ such that for any two probability distributions $\mu,\nu$ supported on the unit ball we have
$\mathcal{E}_{\gamma}(\mu,\nu)\leq c\mathsf{TV}(\mu,\nu)\,.$
Proof.
Assume first that both $\mu,\nu$ are absolutely continuous, and let $f(x)=\frac{\mathrm{d}\mu}{\mathrm{d}x}-\frac{\mathrm{d}\nu}{\mathrm{d}x}$ and $\epsilon=\mathsf{TV}(\mu,\nu)$. By 2.8, it suffices to upper bound $\|I_{s}f\|_{2}$ for $s=(d+\gamma)/2$. First we decompose $\|I_{s}f\|_{2}$ as
$\|I_{s}f\|_{2}\leq\|I_{s}f\mathbbm{1}_{\mathbb{B}(0,2)^{c}}\|_{2}+\|I_{s}f\mathbbm{1}_{\mathbb{B}(0,2)}\|_{2}$
by the triangle inequality. Let $f^{+}(x)=\max\{f(x),0\}$ and $f^{-}(x)=\max\{-f(x),0\}$ so that $\int f^{+}(x)\mathrm{d}{x}=\int f^{-}(x)\mathrm{d}{x}=\epsilon$. Since $\operatorname{supp}(f)\subseteq\mathbb{B}(0,1)$, for all $\|x\|>2$ we have
$c_{s}I_{s}f(x)$ $=\int\frac{f^{+}(y)-f^{-}(y)}{\|x-y\|^{d-s}}\mathrm{d}{y}$
$\leq{{\int f^{+}(y)\mathrm{d}{y}}\over(\|x\|-1)^{d-s}}-{\int f^{-}(y)\mathrm{d}{y}\over(\|x\|+1)^{d-s}}$
$=\epsilon\left({1\over(\|x\|-1)^{d-s}}-{1\over(\|x\|+1)^{d-s}}\right)$
$\leq\epsilon\frac{2(d-s)}{(\|x\|-1)^{d-s+1}}$
[…]
$=\frac{2(d-s)\sqrt{\sigma(\mathbb{S}^{d-1})}\epsilon}{c_{s}}\left(\int_{2}^{\infty}\frac{u^{d-1}}{(u-1)^{2d-2s+2}}\mathrm{d}{u}\right)^{1/2}\lesssim\epsilon$
where we discard a finite constant depending only on $d$ and $\gamma$.
Next we need to estimate $\|I_{s}f\mathbbm{1}_{\mathbb{B}(0,2)}\|_{2}$. Let $q={2\over 1-\gamma/d}>2$ and let $\|\cdot\|_{q,w}$ denote the weak $q$-norm. Define the distribution function $\lambda(t)=m\{x\in\mathbb{R}^{d}:|I_{s}f(x)\mathbbm{1}_{\mathbb{B}(0,2)}|>t\}$ where $m$ is the Lebesgue measure on $\mathbb{R}^{d}$. Because $\lambda(t)\leq\min\left(t^{-q}\|I_{s}f\mathbbm{1}_{\mathbb{B}(0,2)}\|_{q,w}^{q},m(\mathbb{B}(0,2))\right)$ for any $t\geq 0$, we have
$\|I_{s}f\mathbbm{1}_{\mathbb{B}(0,2)}\|_{2}=\left(2\int_{0}^{\infty}t\lambda(t)\mathrm{d}{t}\right)^{\frac{1}{2}}\leq C_{2}\|I_{s}f\mathbbm{1}_{\mathbb{B}(0,2)}\|_{q,w}$
where $C_{2}(d,\gamma)=(\frac{q}{q-2})^{\frac{1}{2}}m(\mathbb{B}(0,2))^{\frac{1}{2}-\frac{1}{q}}$. By the Hardy-Littlewood-Sobolev lemma [Ste70, Theorem V.1],
$\|I_{s}f\|_{q,w}\lesssim\|f\|_{1}=2\epsilon,$
where we again discard a constant depending on $d$ and $\gamma$. Combining these inequalities together, we get
$\|I_{s}f\|_{2}\leq\|I_{s}f\mathbbm{1}_{\mathbb{B}(0,2)^{c}}\|_{2}+\|I_{s}f\mathbbm{1}_{\mathbb{B}(0,2)}\|_{2}\leq C(d,\gamma)\epsilon$
for some constant $C(d,\gamma)$.
Finally, if either $\mu$ or $\nu$ does not have a density, we pick a positive mollifier $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$ such that $\text{supp}(\phi)\subseteq\mathbb{B}(0,1)$, $\phi\geq 0$, and $\int\phi=1$ and set $\phi_{\eta}(x)=\eta^{-d}\phi(x-\eta)$ for $\eta>0$. Consider $\mu_{\eta}=\mu*\phi_{\eta}$ and $\nu_{\eta}=\nu*\phi_{\eta}$, which are both absolutely continuous and supported on $\mathbb{B}(0,1+\eta)$. By the first part of our proof so far we get $\mathcal{E}_{\gamma}(\mu_{\eta},\nu_{\eta})\leq C(d,\gamma)(1+\eta)^{\gamma/2}\mathsf{TV}(\mu_{\eta},\nu_{\eta})$ by a simple argument that rescales the two mollified distributions to be supported on the unit ball. It is well known that $\mu_{\eta}\to\mu$ and $\nu_{\eta}\to\nu$ as $\eta\to 0$ in a weak sense, thus by the definition of $\mathcal{E}_{\gamma}(\mu_{\eta},\nu_{\eta})$ given in (2.1), we obtain $\mathcal{E}_{\gamma}(\mu_{\eta},\nu_{\eta})\to\mathcal{E}_{\gamma}(\mu,\nu)$ as $\eta\to 0$. Moreover, we have $\mathsf{TV}(\mu_{\eta},\nu_{\eta})\leq\mathsf{TV}(\mu,\nu)$ by the data processing inequality, which concludes the proof.
3.2 Lower bound – smooth distributions and Gaussian mixtures
In Section 3.1 we showed that the energy distance is upper bounded by total variation for compactly supported measures. In this section we look at the reverse direction, namely, we aim to lower bound the energy distance by total variation.
Theorem 3.
For any $\beta>0$, $d\geq 1$ and $C>0$, there exists a finite constant $C_{1}$ so that
$\sqrt{\gamma(2-\gamma)}\mathsf{TV}(\mu,\nu)^{\frac{2\beta+d+\gamma}{2\beta}}\leq C_{1}\mathcal{E}_{\gamma}(\mu,\nu)$ (3.1)
for any $\mu,\nu\in\mathcal{P}_{S}(\beta,d,C)$ and $\gamma\in(0,2)$. Similarly, for any $d\geq 1$ there exists a finite constant $C_{2}$ such that
$\frac{\sqrt{\gamma(2-\gamma)}\mathsf{TV}(\mu,\nu)}{\log(3+1/\mathsf{TV}(\mu,\nu))^{\frac{2d+\gamma}{4}}}\leq C_{2}\mathcal{E}_{\gamma}(\mu,\nu)$ (3.2)
for every $\mu,\nu\in\mathcal{P}_{G}(d)$ and $\gamma\in(0,2)$.
Proof.
Abusing notation, identify $\mu$ and $\nu$ with their Lebesgue densities. The argument proceeds trough a chain of inequalities:
  1. 1.
    Bound $\mathsf{TV}$ by the $L^{2}$ distance between densities.
  2. 2.
    Apply Parseval’s Theorem to pass to Fourier space.
  3. 3.
    Apply Hölder’s inequality with well-chosen exponents.
Proof of (3.1). Jensen’s inequality implies that
$2\mathsf{TV}(\mu,\nu)=\|\mu-\nu\|_{1}\leq\sqrt{\operatorname{vol}(\mathbb{B}(0,1))}\|\mu-\nu\|_{2}\lesssim\|\mu-\nu\|_{2},$
where $\operatorname{vol}$ denotes volume and we discard dimension-dependent constants. This completes the first step of our proof. For the second step note that $\mu,\nu\in L^{2}(\mathbb{R}^{d})$ and we may apply Parseval’s theorem to obtain
$\|\mu-\nu\|_{2}^{2}=\frac{1}{(2\pi)^{d}}\|\widehat{\mu}-\widehat{\nu}\|_{2}^{2}.$
For arbitrary $\varphi>0$ and $r\in[1,\infty]$, Hölder’s inequality with exponents $\frac{1}{r}+\frac{1}{r^{*}}=1$ implies that
$\|\widehat{\mu}-\widehat{\nu}\|_{2}^{2}$ $=\int_{\mathbb{R}^{d}}|\widehat{\mu}(\omega)-\widehat{\mu}(\omega)|^{2}\frac{\|\omega\|^{\varphi}}{\|\omega\|^{\varphi}}\mathrm{d}\omega$ (3.3)
$\leq\left(\int_{\mathbb{R}^{d}}|\widehat{\mu}(\omega)-\widehat{\nu}(\omega)|^{2}\|\omega\|^{\varphi r}\mathrm{d}\omega\right)^{1/r}\left(\int_{\mathbb{R}^{d}}\frac{|\widehat{\mu}(\omega)-\widehat{\nu}(\omega)|^{2}}{\|\omega\|^{\varphi r^{*}}}\mathrm{d}\omega\right)^{1/r^{*}}.$
Now, we choose $\varphi$ and $r$ to satisfy
$\varphi r$ $=2\beta$
$\varphi r^{*}$ $=d+\gamma.$
The first equation ensures that the first integral term is bounded by $\|\mu-\nu\|_{\beta,2}^{2/r}$, which is assumed to be at most a $d,\beta$ dependent constant. The second equation ensures that the second integral term is equal to $(\mathcal{E}_{\gamma}(\mu,\nu)^{2}/F_{\gamma}(d))^{1/r^{*}}$ by Proposition 3. The solution to this system of equations is given by $r^{*}=(2\beta+d+\gamma)/(2\beta)$ and $\varphi=2\beta\cdot\frac{d+\gamma}{2\beta+d+\gamma}$. Note that clearly $\varphi>0$ and $r^{*}\geq 1$. Thus, after rearrangement and using that $F_{d}(\gamma)=\Theta(\gamma(2-\gamma))$ up to a dimension dependent constant, we obtain
$\sqrt{\gamma(2-\gamma)}\|\widehat{\mu}-\widehat{\nu}\|_{2}^{\frac{2\beta+d+\gamma}{2\beta}}\leq C_{1}\mathcal{E}_{\gamma}(\mu,\nu),$
for a finite constant $C_{1}=C_{1}(d,\beta)$, concluding the proof.
Proof of (3.2). We write $C(d)\in(0,\infty)$ for a dimension dependent constant that may change from line to line. The outline of the argument is analogous to the above, with the additional step of having to bound the $(\beta,2)$-Sobolev norm of the Gaussian density as $\beta\to\infty$ for which we rely on Lemma 8. Let $\mu$ and $\nu$ have densities $p*\phi$ and $q*\phi$, where $\phi$ is the density of $\mathcal{N}(0,I_{d})$. Writing $f=(p-q)*\phi$, we can extend the proof of [JPW23, Theorem 22] to multiple dimensions to find, for any $R>2$, that
$2\mathsf{TV}(\mu,\nu)=\|\mu-\nu\|_{1}$ $=\int_{\|x\|\leq R}\left|(f*\phi)(x)\right|\mathrm{d}x+\int_{\|x||>R}\left|\int_{\mathbb{R}^{d}}\phi(x-y)\mathrm{d}f(y)\right|\mathrm{d}x$
$\leq\sqrt{\operatorname{vol}_{d}(\mathbb{B}(0,R))}\sqrt{\int_{\|x\|\leq R}|(f*\phi)(x)|^{2}\mathrm{d}x}+\int_{\|x\|>R}\exp(-\|x\|^{2}/8)\mathrm{d}x$
$\leq C(d)\Big{(}R^{d/2}\|\mu-\nu\|_{2}+\exp(-\Omega(R^{2}))\Big{)},$
where the second line uses that $\operatorname{supp}(f)\subseteq\mathbb{B}(0,1)$. Taking $R\asymp\sqrt{\log(3+1/\|\mu-\nu\|_{2})}$ we obtain the inequality
$\mathsf{TV}(\mu,\nu)\leq C(d)\|\mu-\nu\|_{2}\log(3+1/\|\mu-\nu\|_{2})^{d/4}.$ (3.4)
By Hölder’s inequality we obtain
$\|\widehat{f}\|_{2}$ $\leq\|\|\omega\|^{\beta}\widehat{f}(\omega)\|_{2}^{\frac{d+\gamma}{2\beta+d+\gamma}}\left\|\frac{\widehat{f}(\omega)}{\|\omega\|^{\frac{d+\gamma}{2}}}\right\|_{2}^{\frac{2\beta}{2\beta+d+\gamma}}$
$=\|\|\omega\|^{\beta}\widehat{f}(\omega)\|_{2}^{{}^{\frac{d+\gamma}{2\beta+d+\gamma}}}\cdot\mathcal{E}_{\gamma}(\mu,\nu)^{{\frac{2\beta}{2\beta+d+\gamma}}}\cdot F_{\gamma}(d)^{-\frac{\beta}{2\beta+d+\gamma}}$
by Proposition 3. Using that $|\widehat{f}|\leq|\widehat{\phi}|$ and applying Lemma 8, for $\beta\geq 1$ we get
$F_{\gamma}(d)^{\frac{\beta}{2\beta+d+\gamma}}\|\widehat{f}\|_{2}\leq\mathcal{E}_{\gamma}(\mu,\nu)^{{\frac{2\beta}{2\beta+d+\gamma}}}\left(\frac{5\pi^{d/2}}{\Gamma(d/2)}\left(\frac{2\beta+d}{2e}\right)^{\frac{2\beta+d-1}{2}}\right)^{\frac{d+\gamma}{2(2\beta+d+\gamma)}}.$
Rearranging and using Parseval’s Theorem, we get
$\mathcal{E}_{\gamma}(\mu,\nu)\geq C(d)\sqrt{\gamma(2-\gamma)}\|f\|_{2}\frac{\|f\|_{2}^{\frac{d+\gamma}{2\beta}}}{\left(\frac{2\beta+d}{2e}\right)^{\frac{(d+\gamma)(2\beta+d-1)}{8\beta}}}$
for some $d$-dependent, albeit exponential, constant $C(d)>0$. Plugging in $\beta=\log(3+1/\|f\|_{2})$ and assuming that $\|f\|_{2}$ is small enough in terms of $d$, we obtain
$\mathcal{E}_{\gamma}(\mu,\nu)$ $\geq\frac{C(d)\sqrt{\gamma(2-\gamma)}\|f\|_{2}}{\log(3+1/\|f\|_{2})^{\frac{d+\gamma}{4}}}\geq\frac{C(d)\sqrt{\gamma(2-\gamma)}\mathsf{TV}(\mu,\nu)}{\log(3+1/\mathsf{TV}(\mu,\nu))^{\frac{2d+\gamma}{4}}},$ (3.5)
where the second inequality uses (3.4) and Lemma 4. ∎
Theorem 3 is our main technical result, which shows that $\mathcal{E}_{\gamma}$ is lower bounded by a polynomial of the total variation distance for both the smooth distribution class $\mathcal{P}_{S}$ and Gaussian mixtures $\mathcal{P}_{G}$. Note also that in one dimension, (3.1) follows from the Gagliardo–Nirenberg-Sobolev interpolation inequality. However, to our knowledge, the inequality is new for $d>1$. As for the tightness of Theorem 3, we manage to prove that this inequality is the best possible for $\mathcal{P}_{S}$ in one dimension, and best possible up to a poly-logarithmic factor in dimension $2$ and above.
Proposition 7.
For any $\beta>0$, $d\geq 1$, $\gamma\in(0,2)$ and $C>0$ satisfying 1, there exists a finite constant $C_{1}$ so that for any value of $\epsilon\in(0,1)$, there exist $\mu_{\epsilon},\nu_{\epsilon}\in\mathcal{P}_{S}(\beta,d,C)$ such that $\mathsf{TV}(\mu_{\epsilon},\nu_{\epsilon})/\epsilon\in(1/C_{1},C_{1})$ and
$\mathcal{E}_{\gamma}(\mu_{\epsilon},\nu_{\epsilon})\leq C_{1}\mathsf{TV}(\mu_{\epsilon},\nu_{\epsilon})^{\frac{2\beta+d+\gamma}{2\beta}}\log\left(3+\frac{1}{\mathsf{TV}(\mu_{\epsilon},\nu_{\epsilon})}\right)^{d-1}.$
In the special case $\gamma=1$ we obtain an even stronger notion of tightness.
Proposition 8.
When $\gamma=1$ we may replace $\mathcal{E}_{1}$ by $\overline{d_{H}}$ in Proposition 7.
Proposition 8 is an improvement over Proposition 7 due to the inequality $d_{H}\lesssim\overline{d_{H}}$ over the class $\mathcal{P}_{S}(\beta,d,C)$, which follows from LABEL:prop:d_H<_d_H. It shows also that our construction has the property that there does not exist any halfspace that separates $\mu$ and $\nu$ better than our bounds suggest.
The proofs of both results are presented in Appendix D. The general idea is to saturate Hölder’s inequality in (3.3), for which the Fourier transform of $f=\frac{\mathrm{d}\nu}{\mathrm{d}x}-\frac{\mathrm{d}\mu}{\mathrm{d}x}$ should be supported on a sphere. However, such $f$ clearly cannot be compactly supported. Thus the actual construction is to multiply the Fourier inverse of the uniform measure on a sphere with a compactly supported mollifier. In $d>1$ the mollifier that we require must have super-polynomial Fourier spectrum decay, for which we use the recent construction in [Coh23].
3.3 Lower bound – discrete distributions
Suppose we have two discrete distributions that are supported on a common, finite set of size $k$. One way to measure the energy distance between them would be to identify their support with the set $\{1,2,\dots,k\}$, thereby embedding the two distributions in $\mathbb{R}$, and applying the one-dimensional energy distance.
While the above approach seems reasonable, it is entirely arbitrary. Indeed, there might not be a natural ordering of the support; moreover, why should one choose the integers between $1$ and $k$ instead of, say, the set $\{1,2,4,\dots,2^{k}\}$? The total variation distance doesn’t suffer from such ambiguities, and it is unclear how our choice of embedding affects the relationship to $\mathsf{TV}$. The following result attacks precisely this question.
Theorem 4.
Let $\mu$ and $\nu$ be probability distributions supported on the set $\{x_{1},\dots,x_{k}\}\subseteq\mathbb{R}^{d}$ and let $\delta=\min_{i\neq j}\|x_{i}-x_{j}\|$. Then there exists a universal constant $C>0$ such that
$\mathcal{E}_{1}^{2}(\mu,\nu)\geq\frac{C\delta}{k\sqrt{d}}\mathsf{TV}^{2}(\mu,\nu).$
Proof.
Let $\mu=\sum_{i=1}^{k}\mu_{i}\delta_{x_{i}}$ and $\nu=\sum_{i=1}^{k}\nu_{i}\delta_{x_{i}}$. Then, by [Bal92, Theorem 1] we have
$\mathcal{E}^{2}_{1}(\mu,\nu)$ $=-\sum_{i,j}(\mu_{i}-\nu_{i})(\mu_{j}-\nu_{j})\|x_{i}-x_{j}\|\geq\frac{C\delta}{\sqrt{d}}\sum_{i=1}^{k}(\mu_{i}-\nu_{i})^{2}\geq\frac{C\delta\mathsf{TV}^{2}(\mu,\nu)}{k\sqrt{d}}$
as required.
Remark 2.
Similar results can be proved for the generalized energy distance $\mathcal{E}_{\gamma}$, using e.g. the work [NW92]. However, to the best of our knowledge, these estimates degrade significantly in the dimension $d$ in contrast with Ball’s result [Bal92].
Notice that by our discussion above, the support set $\{x_{1},\dots,x_{k}\}$ in Theorem 4 is arbitrary and may be chosen by us. Since the scale of the supporting points $x_{1},\dots,x_{k}$ is statistically irrelevant, we remove this ambiguity by restricting the points to lie in the unit ball, i.e. requiring that $\max_{i}\|x_{i}\|\leq 1$. We see now that the comparison between $\mathcal{E}_{1}$ and $\mathsf{TV}$ improves as $\delta/\sqrt{d}$ grows. Given a fixed value of $\delta$, we want to make the dimension $d$ of our embedding as low as possible, which means that the points $x_{1},\dots,x_{k}$ should form a large $\delta$-packing of the $d$-dimensional unit ball. Due to well known bounds on the packing number of the Euclidean ball, it follows that the best one can hope for is
$\log(k)\asymp d\log(1/\delta).$
Maximizing $\delta/\sqrt{d}$ subject to this constraint yields the choice $d=\Theta(\log(k))$ and $\delta=\Theta(1)$. This gives us the following corollary.
Corollary 5.
There exists a universal constant $C\in(0,\infty)$ such that for any $k\geq 1$ there exists a set of points $x_{1},\dots,x_{k}\in\mathbb{R}^{\lceil{C\log(k)}\rceil}$ with $\max_{i}\|x_{i}\|\leq 1$ such that
$\mathcal{E}_{1}\left(\sum_{i=1}^{k}\mu_{i}\delta_{x_{i}},\sum_{i=1}^{k}\nu_{i}\delta_{x_{i}}\right)\geq\frac{\mathsf{TV}(\mu,\nu)}{C\sqrt{k}\sqrt[4]{\log(k)}}$
for any two probability mass functions $\mu=(\mu_{1},\dots,\mu_{k})$ and $\nu=(\nu_{1},\dots,\nu_{k})$.
The question arises how the set of points $x_{1},\dots,x_{k}$ in Corollary 5 should be constructed. One solution is to use an error correcting code (ECC), whereby we take the $x_{i}$ to be the codewords of an ECC on the scaled hypercube $\frac{1}{\sqrt{d}}\{\pm 1\}^{d}$ for some dimension $d$ (known as “blocklength” in this context). An ECC is asymptotically good if the message length $\log(k)$ is linear in the blocklength $d$, that is $d\asymp\log(k)$, and if the minimum Hamming distance between any two codewords is $\Theta(d)$, which translates precisely into $\delta=\min_{i\neq j}\|x_{i}-x_{j}\|\asymp 1$. Many explicit constructions of asymptotically good error correcting codes exist, see [Jus72] for one such example, and random codes are almost surely good [BF02]. Clearly the better the code is, the better the constants we obtain in Corollary 5.
Remark 3.
One interesting consequence of Corollary 5 and the preceeding discussion is the following: given a categorical feature with $k$ possible values, the perceptron may obtain better performance by identifying each category with the codewords $x_{1},\dots,x_{k}$ of an ECC instead of the standard one-hot encoding.
4 Density estimation
In this section we apply what we’ve learnt about the generalized energy distance and the perceptron discrepancy in prior sections, and analyze multiple problems related to density estimation.
4.1 Estimating smooth distributions and Gaussian mixtures
Suppose that $X_{1},\dots,X_{n}\stackrel{{\scriptstyle iid}}{{\sim}}\nu$ for some probability distribution $\nu$ on $\mathbb{R}^{d}$. Given a class of “generator” distributions $\mathcal{G}$ and $\gamma\in(0,2)$, define the minimum-$\mathcal{E}_{\gamma}$ estimator as
$\tilde{\nu}_{\gamma}\in\operatorname*{arg\,min}_{\nu^{\prime}\in\mathcal{G}}\mathcal{E}_{\gamma}(\nu^{\prime},\nu_{n}),$ (4.1)
where $\nu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$. Note that $\tilde{\nu}_{\gamma}$ doesn’t quite agree with our definition of $\tilde{\nu}_{\gamma}$ in (1.2), because the $\gamma=1$ case minimizes the average halfspace distance $d_{H}\asymp\mathcal{E}_{1}$ and not the perceptron discrepancy $\overline{d_{H}}$. The following two results bound the performance of $\tilde{\nu}$ as defined in (4.1), as an estimator of $\nu$ for the smooth density class $\mathcal{P}_{S}$ as well as the Gaussian mixture class $\mathcal{P}_{G}$. In Section 4.2 we present the adapation of these to $\overline{d_{H}}$, thereby proving Theorem 1.
Theorem 6.
Let $\tilde{\nu}_{\gamma}$ be the estimator defined in (4.1). For any $\beta>0$, $d\geq 1$ and $C>0$, there exists a finite constant $C_{1}$ so that
$\sup_{\nu\in\mathcal{P}_{S}(\beta,d,C)}\mathbb{E}\mathsf{TV}(\tilde{\nu}_{\gamma},\nu)\leq C_{1}(n\gamma(2-\gamma))^{-\frac{\beta}{2\beta+d+\gamma}}$ (4.2)
holds for $\mathcal{G}=\mathcal{P}_{S}(\beta,d,C)$ and any $\gamma\in(0,2)$. Similarly, for any $d\geq 1$ there is a finite constant $C_{2}$ such that
$\sup_{\nu\in\mathcal{P}_{G}(d)}\mathbb{E}\mathsf{TV}(\tilde{\nu}_{\gamma},\nu)\leq C_{2}\phi(n\gamma(2-\gamma))$ (4.3)
holds for $\mathcal{G}=\mathcal{P}_{G}(d)$ and any $\gamma\in(0,2)$, where $\phi(x)=\frac{\log(3+x)^{\frac{2d+\gamma}{4}}}{\sqrt{x}}$.
Proof.
Let us focus on the case $\mathcal{G}=\mathcal{P}_{S}(\beta,d,C)$ first and let $t=\frac{2\beta+d+\gamma}{2\beta}$. The inequality $\mathcal{E}_{\gamma}(\tilde{\nu}_{\gamma},\nu_{n})\leq\mathcal{E}_{\gamma}(\nu,\nu_{n})$ holds almost surely by the definition of $\tilde{\nu}_{\gamma}$. Writing $C_{1}=C_{1}(\beta,d,C)$ for a finite constant that we relabel freely, the first claim is substantiated by the chain of inequalities
[…]
$\stackrel{{\scriptstyle\Delta-\text{ineq.}}}{{\leq}}\mathbb{E}\Big{[}\left(C_{1}\frac{\mathcal{E}_{\gamma}(\nu,\nu_{n})+\mathcal{E}_{\gamma}(\tilde{\nu}_{\gamma},\nu_{n})}{\sqrt{\gamma(2-\gamma)}}\right)^{1/t}\Big{]}$
[…]
$\,\,\stackrel{{\scriptstyle\text{Jensen}}}{{\leq}}\left(2C_{1}\frac{\mathbb{E}\mathcal{E}_{\gamma}(\nu,\nu_{n})}{\sqrt{\gamma(2-\gamma)}}\right)^{1/t}$
[…]
The result for $\mathcal{G}=\mathcal{P}_{G}$ follows analogously. Define the function $r(x)=x\sqrt{\gamma(2-\gamma)}/\log(3+1/x)^{\frac{2d+\gamma}{4}}$. One can check by direct calculation that $r$ is strictly increasing and convex on $\mathbb{R}_{+}$. As a consequence, its inverse $r^{-1}$ is strictly increasing and concave. Let $C_{2}$ be a $d$-dependent finite constant which we relabel repeatedly. Similarly to the case of smooth distributions covered above, using Theorem 3 and Jensen’s inequality we obtain the chain of inequalities
$\mathbb{E}\mathsf{TV}(\tilde{\nu}_{\gamma},\nu)$ $\leq\mathbb{E}\Big{[}r^{-1}\left(C_{2}\frac{\mathcal{E}_{\gamma}(\tilde{\nu}_{\gamma},\nu)}{\sqrt{\gamma(2-\gamma)}}\right)\Big{]}$
$\leq r^{-1}\left(C_{2}\frac{\mathbb{E}\mathcal{E}_{\gamma}(\tilde{\nu}_{\gamma},\nu)}{\sqrt{\gamma(2-\gamma)}}\right)$
$\leq r^{-1}\left(2C_{2}\frac{\mathbb{E}\mathcal{E}_{\gamma}(\nu,\nu_{n})}{\sqrt{\gamma(2-\gamma)}}\right)$
$\leq r^{-1}\left(C_{2}(n\gamma(2-\gamma))^{-1/2}\right).$
The conclusion follows by Lemma 4. ∎
Notice that the rate of estimation of the minimum $\mathcal{E}_{\gamma}$ density estimator improves as $\gamma\downarrow 0$, and in fact seems to approach the optimum. However, simultaneously, the “effective sample size” $n\gamma$ shrinks. The best trade-off that we can derive is the following.
Corollary 7.
The rate in (4.2) (resp. (4.3)) can be improved to $(\log(n)/n)^{\beta/(2\beta+d)}$ (resp. $\log(n)^{d/4}\sqrt{\log\log n}/\sqrt{n}$) by setting $\gamma=\log(n)^{-1}$ (resp. $\gamma=\log\log(n)^{-1}$) adaptively.
4.2 Proof of Theorem 1 and Theorem 2
We already have everything needed to deduce Theorem 2. Since it is an exercise in combining results, we simply list the required steps:
  1. 1.
    Use Theorem 3 to get a comparison between $\mathsf{TV}$ and $\mathcal{E}_{\gamma}$.
  2. 2.
    Set $\gamma=1$ and use Proposition 2 to get the equivalence between $\mathcal{E}_{1}$ and $d_{H}$.
  3. 3.
    Use Proposition 1 to get a comparison between $d_{H}$ and $\overline{d_{H}}$.
Turning to the proof of Theorem 1, we find that it is completely analogous to the proof of Theorem 6, with the only difference being that we can no longer rely on Lemma 1 to show that the distance between empirical and population measures decays at the parametric rate, as the latter applies to $\mathcal{E}_{\gamma}$ instead of $\overline{d_{H}}$. However, the corresponding result for $\overline{d_{H}}$ is well known.
Lemma 2.
Let $\nu$ be a probability distribution on $\mathbb{R}^{d}$ and $\nu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$ for i.i.d. observations $X_{i}\sim\nu$. Then, for a finite universal constant $C$,
$\mathbb{E}\overline{d_{H}}(\nu,\nu_{n})\leq C\sqrt{\frac{d}{n}}.$
Proof.
Follows by e.g. [Ver18, 8.3.23] and the fact that $\mathcal{D}_{a}$, the family of halfspace indicators, has VC dimension $d+1$. ∎
With Lemma 2 in hand, completing the argument is straightforward: To deduce Theorem 1 follow the same steps as in the proof of Theorem 6, except use Theorem 2 and Lemma 2 in place of Theorem 3 and Lemma 1 respectively.
4.3 Estimating discrete distributions
In many practical machine learning tasks the data is discrete, albeit on a large alphabet $[k]=\{1,2,\dots,k\}$: for example, in recommender systems the alphabet could be all possible ads, products or articles. A common idea to apply modern learning pipelines to such data is to use an embedding $E:[k]\to\mathbb{R}^{d}$, with “one-hot” encoding ($d=k$) being the most popular choice. After such an embedding, the data is effectively made “continuous” and the density estimation methods as discussed previously can be applied. Can such an approach be good in the sense of minimax estimation guarantees? We answer this question positively in this section, provided that embedding $E$ comes from an error-correcting code.
Let $\mathcal{P}_{k}$ denote the set of all probability distributions on the set $[k]$. Suppose we observe an i.i.d. sample $X_{1},\dots,X_{n}\sim\nu$ from some unknown distribution $\nu\in\mathcal{P}_{k}$. The problem of estimating $\nu$ is effectively trivial: the empirical distribution provides a minimax optimal estimator. Indeed, it is a folklore fact (see e.g. [Can20, Theorem 1] or [PW23, Exc. VI.8]) that the optimal rate of estimation is given by
$\sup_{\nu\in\mathcal{P}_{k}}\mathbb{E}\mathsf{TV}^{2}\left(\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}},\nu\right)\asymp\min\left\{\frac{k}{n},1\right\}.$ (4.4)
Recall from Section 3.3 that we may choose to embed the alphabet $[k]$ into some higher dimensional Euclidean space. Given distinct points $x_{1},\dots,x_{k}\in\mathbb{R}^{d}$ for some $d\geq 1$, we can identify any distribution $\mu\in\mathcal{P}_{k}$ with the probability distribution $\sum_{i=1}^{k}\mu_{i}\delta_{x_{i}}$, where $\mu_{i}$ is the mass that $\mu$ puts on $i\in[k]$.
Theorem 8.
There exists a universal constant $C<\infty$ with the following property. For any alphabet size $k$ there exist embedding points $a_{1},\dots,a_{k}\in\mathbb{R}^{\lceil C\log(k)\rceil}$ such that given an i.i.d. sample $X_{1},\dots,X_{n}\sim\nu$ from an unknown $\nu\in\mathcal{P}_{k}$, any estimator $\tilde{\nu}\in\mathcal{P}_{k}$ that satisfies
$\mathcal{E}_{1}^{2}\left(\sum_{i=1}^{k}\tilde{\nu}_{i}\delta_{a_{i}},\frac{1}{n}\sum_{i=1}^{n}\delta_{a_{X_{i}}}\right)\leq\frac{c\sqrt{d}}{n}$ (4.5)
enjoys the performance guarantee
$\sup_{\nu\in\mathcal{P}_{k}}\mathbb{E}\mathsf{TV}^{2}(\tilde{\nu},\nu)\leq C\min\left\{(c+1)\frac{k\log(k)}{n},1\right\}.$ (4.6)
Moreover, we may replace $\mathcal{E}_{1}$ by $\overline{d_{H}}$ in (4.5) and the result (4.6) remains true with an additional multiplicative $\sqrt{\log k}$ factor.
Proof.
Let $a_{1},\dots,a_{k}\in\mathbb{R}^{d}$ be the points defined in Corollary 5 (relabeled from $x_{1},\dots,x_{k}$ for clarity) so that $d\asymp\log(k)$. By the triangle inequality we have
$\mathbb{E}\mathcal{E}^{2}_{1}\left(\sum_{i=1}^{k}\tilde{\nu}_{i}\delta_{a_{i}},\sum_{i=1}^{k}\nu_{i}\delta_{a_{i}}\right)$ $\leq 2\mathbb{E}\mathcal{E}^{2}_{1}\left(\sum_{i=1}^{k}\tilde{\nu}_{i}\delta_{a_{i}},\frac{1}{n}\sum_{i=1}^{n}\delta_{a_{X_{i}}}\right)+2\mathbb{E}\mathcal{E}^{2}_{1}\left(\frac{1}{n}\sum_{i=1}^{n}\delta_{a_{X_{i}}},\sum_{i=1}^{k}\nu_{i}\delta_{a_{i}}\right)$
[…]
By Corollary 5, the definition of $\tilde{\nu}$ and the triangle inequality it follows that
$\mathbb{E}\mathsf{TV}^{2}(\tilde{\nu},\nu)\lesssim\frac{kd(c+1)}{n}\asymp\frac{k\log(k)(c+1)}{n}.$
Noting the trivial fact that $\mathsf{TV}\leq 1$ completes the proof of the first claim.
Suppose now that we replace $\mathcal{E}_{1}$ by $\overline{d_{H}}$ in the definition of $\tilde{\nu}$. The proof follows analogously, using the chain of inequalities
[…]
and Lemma 2 in place of Lemma 1, which is where we loose the $\sqrt{d}\asymp\sqrt{\log(k)}$ factor. ∎
As we explained, the empirical distribution $\tilde{\nu}=\frac{1}{n}\sum_{i=1}\delta_{X_{i}}$ achieves optimality in (4.4), and clearly also achieves $c=0$ in (4.5) i.e. minimizes the empirical risk globally. The point of Theorem 8 is to show that approximate minimizers, such as those found via SGD, are also nearly minimax optimal.
4.3.1 Estimating Hölder smooth densities
Theorem 8 has interesting implications for density estimation over the class of distributions on the cube $[0,1]^{d}$ with uniformly bounded derivatives up to order $\underline{\beta}\triangleq\lceil\beta-1\rceil$ and $(\beta-\underline{\beta})$-Hölder continuous $\underline{\beta}^{th}$ derivative; call such distributions simply $\beta$-Hölder smooth [^5] Note that this class is not the same as $\mathcal{P}_{S}$, although related.. Writing $B_{j}$ for the cube with center $(j-\frac{1}{2})\epsilon^{1/\beta}$ and sidelength $\epsilon^{1/\beta}$ where $j\in\{1,\dots,\epsilon^{-1/\beta}\}^{d}$, it is known (see e.g. [ACPS18, Lemma 7.2] or [IS03, Proposition 2.16]) that
$\sum_{j}\left|\int_{B_{j}}(f(x)-g(x))\mathrm{d}x\right|=\int_{[0,1]^{d}}|f(x)-g(x)|\mathrm{d}x+O(\epsilon)$
for any $\beta$-Hölder smooth densities $f,g$. In other words, discretizing such distributions using a regular grid with $\Omega(\epsilon^{-d/\beta})$ cells maintains total-variation distances up to an additive $O(\epsilon)$ error.
Now, consider a ‘multilayer perceptron’, i.e. a fully connected multilayer neural network with activations given by $x\mapsto\mathbbm{1}\{x\geq 0\}$. Such a multilayer network with large enough hidden layers can in principle implement the discretization described above, and embed the $\epsilon^{-d/\beta}$ cells as an error correcting code. Thus, due to Theorem 8, the ERM density estimator (1.2) would achieve the minimax optimal density estimation rate $n^{-\beta/(2\beta+d)}$ over $\beta$-Hölder smooth densities up to polylog factors provided the discriminator class $\mathcal{D}$ includes the aforementioned multilayer perceptron and has VC-dimension at most polylog in $1/\epsilon$. This observation essentially generalizes Theorem 1, which shows that if the discriminator class includes only the single layer perceptron then the best possible minimax rate is $n^{-\beta/(2\beta+d+1)}$.
4.4 A stopping criterion for smooth density estimation
As a corollary to our results, we propose a stopping criterion for training density estimators. Before doing so, let us record a result about the concentration properties of the empirical energy distance about its expectation.
Lemma 3.
Let $\nu$ be supported on a compact subset $\Omega\subseteq\mathbb{R}^{d}$, and let $\nu_{n}$ be its empirical measure based on $n$ i.i.d. observations. For every $\gamma\in(0,2)$ there exists a constant $C_{1}\in(0,\infty)$ such that
$\mathbb{P}\left(\mathcal{E}_{\gamma}(\nu,\nu_{n})\geq\frac{C_{1}}{\sqrt{n}}+t\right)\leq 2\exp\left(-\frac{nt^{2}}{C_{1}}\right).$
In other words, $\mathcal{E}_{\gamma}(\nu,\nu_{n})$ is $O(1/n)$-sub-Gaussian.
Proof.
Recall the MMD formulation of the generalized energy distance from Section 2.3. The corresponding kernel is given by $k_{\gamma}(x,y)=\|x\|^{\gamma}+\|y\|^{\gamma}-\|x-y\|^{\gamma}$. Clearly
$\sup_{x,x^{\prime},y,y^{\prime}\in\operatorname{supp}(\nu)}(k_{\gamma}(x,y)-k_{\gamma}(x^{\prime},y^{\prime}))\lesssim\operatorname{diam}(\Omega)^{\gamma}.$
Therefore, by McDiarmid’s inequality we know that $\mathcal{E}_{\gamma}(\nu,\nu_{n})$ is $O(1/n)$-subGaussian (note we don’t track constants depending on $\Omega$ here). From Lemma 1 we know that $\mathbb{E}\mathcal{E}_{\gamma}(\nu,\nu_{n})\lesssim 1/\sqrt{n}$, and the conclusion follows. ∎
Consider the following scenario: we have i.i.d. training data $X_{1},\dots,X_{n}$ from some distribution $\nu$ and we are training an arbitrary generative model to estimate $\nu$. Suppose that this training process gives us a sequence of density estimators $\{\mu_{k}\}_{k\geq 1}$, which could be the result of, say, subsequent gradient descent steps on our parametric class of generators. Is there any way to figure out after how many steps $K$ we may stop the training process? In other words, can we identify a value of $K$ such that $\mathsf{TV}(\nu,\mu_{K})$ is guaranteed to be less than some threshold with probability $1-\delta$? Note that an additional difficulty here is that our generative model for $\mu_{k}$ is able to generate the samples from $\mu_{k}$ but otherwise gives us no other access to $\mu_{k}$. The fast (dimension-free) concentration properties of $\mathcal{E}_{\gamma}$ and the minimax optimality guarantees of its minimizer (whenever $\nu$ is smooth) make it an excellent choice for such a stopping criteria.
Let $\nu\in\mathcal{P}_{S}(\beta,d,C)$ and let $\nu_{n}$ be its empirical version based on the $n$ i.i.d. observations. Assume further that $\{\mu_{k}\}_{k\geq 1}\subseteq\mathcal{P}_{S}(\beta,d,C)$ is a sequence of density estimators based on the sample $X_{1},\dots,X_{n}$. Finally, given the training sample $(X_{1},\dots,X_{n})$, for each $k$ let $\mu_{k,m_{k}}=\frac{1}{m_{k}}\sum_{i=1}^{m_{k}}\delta_{X^{(k)}_{i}}$ be the empirical distribution of the sample $(X^{(k)}_{1},\dots,X^{(k)}_{m_{k}})\sim\mu_{k}^{\otimes m_{k}}$.
Proposition 9.
For any $\beta>0,d\geq 1$ and $\gamma\in(0,2)$ there exists a constant $c\in(0,\infty)$ such that
$\mathbb{P}\left(\mathsf{TV}(\mu_{k},\nu)\leq c\left(\sqrt{\frac{\log(1/\delta)}{n}}+\mathcal{E}_{\gamma}(\mu_{k,m_{k}},\nu_{n})\right)^{\frac{2\beta}{2\beta+d+\gamma}},\,\,\forall\,k\geq 1\right)\geq 1-2\delta$
provided we take $m_{k}=cn\log(k^{2}/\delta)/\log(1/\delta)$.
Proof.
Let $c=C_{1}$ where $C_{1}$ is as in Lemma 3 and fix $\delta\in(0,1)$. Define the event $A=\left\{\mathcal{E}_{\gamma}(\nu,\nu_{n})\geq\frac{c}{\sqrt{n}}+\sqrt{\frac{c\log(2/\delta)}{n}}\right\}$ and similarly
$A_{k}=\left\{\mathcal{E}_{\gamma}(\mu_{k,m_{k}},\mu_{k})\geq\frac{c}{\sqrt{m_{k}}}+\sqrt{\frac{ct_{k}}{m_{k}}}\right\}$
for some sequence $t_{1},t_{2},\dots,$ and each $k\geq 1$. By Lemma 3,
$\mathbb{P}(A)$ $\leq\delta,$
$\mathbb{P}(A_{k})$ $=\mathbb{E}\mathbb{P}(A_{k}|X_{1},\dots,X_{n})\leq 2\exp(-t_{k}).$
Taking $t_{k}=\log(k^{2}\pi^{2}/(3\delta))$, the union bound gives
$\mathbb{P}\left(A\cup\bigcup_{k\geq 1}A_{k}\right)\leq 2\delta.$
By the inequality $\mathcal{E}_{\gamma}(\mu_{k},\nu)\leq\mathcal{E}_{\gamma}(\mu_{k},\mu_{k,m_{k}})+\mathcal{E}_{\gamma}(\mu_{k,m_{k}},\nu_{n})+\mathcal{E}_{\gamma}(\nu_{n},\nu)$ it follows that
$\mathbb{P}\left(\exists k\geq 1\,:\,\mathcal{E}_{\gamma}(\nu,\mu_{k})>\mathcal{E}_{\gamma}(\mu_{k,m_{k}},\nu_{n})+\frac{c}{\sqrt{n}}+\frac{c}{\sqrt{m}_{k}}+\sqrt{\frac{c\log(2/\delta)}{n}}+\sqrt{\frac{c\log(k^{2}\pi^{2}/(3\delta))}{m_{k}}}\right)\leq 2\delta.$
Thus, by choosing $m_{k}\asymp n\log(k^{2}/\delta)/\log(1/\delta)$ we can conclude that there exists a constant $c^{\prime}$ depending only on $\beta,d,\gamma$ such that
$\mathbb{P}\left(\mathcal{E}_{\gamma}(\nu,\mu_{k})\leq c^{\prime}\sqrt{\frac{\log(1/\delta)}{n}}+\mathcal{E}_{\gamma}(\mu_{k,m_{k}},\nu_{n}),\forall k\geq 1\right)\geq 1-2\delta.$
The final conclusion follows from Theorem 3. ∎
Note that our bound on the probability holds for all $k$ simultaneously, which is made possible by the fact that $m_{k}$ grows as $k\rightarrow\infty$. The empirical relevance of such a result is immediate: suppose we have proposed candidate generative models $\mu_{1},\mu_{2},\ldots$ (e.g. one after each period of training epochs, or from different training models) that is trained on an i.i.d. dataset $X_{1},\dots,X_{n}$ of size $n$ from $\nu\in\mathcal{P}_{S}(\beta,d,C)$. A “verifier” only needs to request for $m_{k}$ independent draws from the $k$’th candidate, and if we ever achieve $\mathcal{E}_{(\log n)^{-1}}\left(\mu_{k,m_{k}},\nu_{n}\right)\lesssim\sqrt{\log(1/\delta)/n}$ we can stop training and claim by Theorem 6 that we are a constant factor away from (near-)minimax optimality with probability $1-\delta$.
5 Suboptimality for two-sample testing
So far in this paper we have shown how the empirical energy distance minimizer, while being mismatched with the target loss (total variation), nevertheless achieves (near) minimax optimal performance for density estimation tasks. Unfortunately, this surprising effect does not carry over to other statistical tasks, such as two-sample testing, which we describe in this section.
The task of two-sample testing over a family of distributions $\mathcal{P}$ is the following. Given two samples $(X,Y)\sim p^{\otimes n}\otimes q^{\otimes m}$ with unknown distribution, we need to distinguish between the hypotheses
$H_{0}:p=q\text{ and }p\in\mathcal{P},\quad\text{versus}\quad H_{1}:\mathsf{TV}(p,q)>\varepsilon,\text{and }p,q\in\mathcal{P}$
with vanishing type-I and type-II error. The special case of $m=\infty$ is known as goodness-of-fit testing, and for the class of smooth distributions it was famously solved by Ingster [Ing87], who showed that in dimension $d=1$ the problem is solvable with probability $1-o(1)$ if and only if
$n=\omega(\epsilon^{-\frac{2\beta+d/2}{\beta}}),$ (5.1)
in which case a variant of the $\chi^{2}$-test works. The case of general $m,n$ and $d\geq 1$ was resolved in [ACPS18] who showed that the problem is also solvable if and only if (5.1) holds with $n$ replaced by $\min\{n,m\}$, using the very same $\chi^{2}$-test; see also [LY19]. In the remainder of the section we focus on the $m=n$ case for simplicity.
In a recent paper [PCGT23], the following test statistic for two-sample testing was proposed:
$T_{d,k}(p,q)=\max_{(w,b)\in\mathbb{S}^{d-1}\times[0,\infty)}\left|\mathbb{E}_{X\sim p}\left(w^{\top}X-b\right)_{+}^{k}-\mathbb{E}_{Y\sim q}\left(w^{\top}Y-b\right)_{+}^{k}\right|$
where the arguments $X,Y$ can be either discrete (e.g. via observed samples) or continuous densities. Note that here we take $(a)^{0}_{+}=\mathbbm{1}\{a\geq 0\}$ by convention. Specifically, the test proposed is to reject the null hypothesis when
$T_{d,k}(p_{n},q_{n})\geq t_{n},$ (5.2)
where $p_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}},q_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{Y_{i}}$ are empirical measures and the threshold that satisfies both $t_{n}=o(1)$ and $t_{n}=\omega(1/\sqrt{n})$. One of their main technical result [PCGT23, Theorem 6] asserts that the test (5.2) returns the correct hypothesis with probability $1-o(1)$ asymptotically as $n\to\infty$ for any qualifying sequence $\{t_{n}\}_{n\geq 1}$ and fixed $p,q$. However, this result leaves open questions about the sample complexity of their test, and in particular, whether it is able to achieve known minimax rates. It turns out that our results imply that their test, at least in the $k=0$ case, cannot attain the optimal two-sample testing sample complexity (5.1) over the smooth class $\mathcal{P}_{S}(\beta,d,C)$. To connect to our results, notice that
$T_{d,0}(p,q)=\overline{d_{H}}(p,q)\,.$
Proposition 10.
For all $d,\beta>0$, there exists constants $c,c^{\prime}$ such that for all $\varepsilon>0$, there exists probability density functions $p,q$ supported on the $d$-dimensional unit ball such that
  1. (i)
    $\|p\|_{\beta,2},\|q\|_{\beta,2}<c$,
  2. (ii)
    $\|p-q\|_{1}\asymp\|p-q\|_{2}\asymp\varepsilon$, and
  3. (iii)
    the expected test statistic satisfies
    $\mathbb{E}[T_{d,0}(p_{n},q_{n})]\leq\frac{c^{\prime}}{\sqrt{n}}\,$
    for any $n\leq(\log{1\over\varepsilon})^{-d}\varepsilon^{-\frac{2\beta+d+1}{\beta}}$.
In other words, consistent testing using the statistic $T_{d,0}$ is impossible with $n=\tilde{o}(\varepsilon^{-\frac{2\beta+d+1}{\beta}})$ samples, which is a far cry from the optimal sample complexity  (5.1) attainable by the $\chi^{2}$ test. The proof of Proposition 10 is given at Appendix E.
6 Conclusion
We analyzed the simple discriminating class of affine classifiers and proved its effectiveness in the ERM-GAN setting (1.2) within the Sobolev class $\mathcal{P}_{S}(\beta,d)$ and Gaussian mixtures $\mathcal{P}_{G}(d)$ with respect to the $L^{2}$ norm (see Theorems 6 and 7) and the total variation distance (see Theorem 1). Our findings affirm the rate’s near-optimality for the considered classes of $\mathcal{P}_{S}$ and $\mathcal{P}_{G}$. Moreover, we present inequalities that interlink the $\mathcal{E}_{\gamma}$, TV, and $L^{2}$ distances, and demonstrate (in some cases) the tightness of these relationships via corresponding lower bound constructions (Appendix D). We also interpret the generalized energy distance in several ways that help advocate for its use in real applications. This work connects to a broader literature on the theoretical analysis of GAN-style models.
An interesting question emerges about the interaction between the expressiveness and concentration of the discriminator class. We found that the class of affine classifiers $\mathcal{D}_{1}$ is guaranteed to maintain some (potentially small) proportion of the total variation distance, and that it decays at the parametric rate between population and empirical distributions. Thus, we have traded off expressiveness for better concentration of the resulting IPM. As discussed in Section 1.2, Yatracos’ estimator lies at the other end of this discriminator expressiveness-concentration trade-off: the distance $d_{Y}$ is as expressive as total variation when restricted to the generator class $\mathcal{G}$, but $\sup_{\nu\in\mathcal{G}}\mathbb{E}d_{Y}(\nu,\nu_{n})$ decays strictly slower than $1/\sqrt{n}$ for nonparametric classes $\mathcal{G}$. A downside compared to $\overline{d_{H}}$ is that $(i)$ the Yatracos class $\mathcal{Y}$ requires knowledge of $\mathcal{G}$ while our $\mathcal{D}_{1}$ is oblivious to $\mathcal{G}$ and $(ii)$ the distance $d_{Y}$ is impractical to compute as it requires a covering of $\mathcal{G}$. Our question is: is it possible to find a class of sets $\mathcal{S}\subseteq 2^{\mathbb{B}(0,1)}$ that lies at an intermediate point on this trade-off? In other words, does $\mathcal{S}$ exist such that the ERM $\tilde{\nu}$ (1.2) using the discriminator class $\mathcal{D}=\mathcal{S}$ is optimal over, say, $\mathcal{G}=\mathcal{P}_{S}$ and the induced distance converges slower than $1/\sqrt{n}$ but faster than $n^{-\beta/(2\beta+d)}$ between empirical and population measures? Would there be desiderata for a sample-efficient discriminator that has neither full expressiveness against total variation and does not concentrate at a parametric rate?
Acknowledgements
Supported in part by the NSF grant CCF-2131115 and the MIT-IBM Watson AI Lab.
Appendix A Auxiliary technical results
In this section we list some technical lemmas that are used in our later proofs.
Theorem 9 (Plancherel theorem).
Let $f,g\in L^{2}(\mathbb{R}^{d})$. Then
$\int_{\mathbb{R}^{d}}f(x)\overline{g(x)}\mathrm{d}x=\frac{1}{(2\pi)^{d}}\int_{\mathbb{R}^{d}}\widehat{f}(\omega)\overline{\widehat{g}(\omega)}\mathrm{d}\omega.$
Lemma 4.
Suppose $t,x,y>0$. Then there exist finite $t$-dependent constants $C_{1},C_{2}$ such that
$x\leq\ y\log(3+1/y)^{t}\implies\frac{x}{\log(3+1/x)^{t}}\leq C_{1}\ y\implies x\leq C_{2}\ y\log(3+1/y)^{t}.$
Proof.
Let us focus on the first implication. If $x\leq y$, then it clearly holds. If $y\leq x\leq y\log(3+1/y)^{t}$ then it suffices to show
$\left(\frac{x}{\log(3+1/x)^{t}}\leq\right)\frac{y\log(3+1/y)^{t}}{\log(3+1/(y\log(3+1/y)^{t}))^{t}}\stackrel{{\scriptstyle!}}{{\leq}}C_{1}y.$
The inequality marked by !!! is equivalent to
$3+1/y\leq(3+1/(y\log(3+1/y)^{t}))^{\sqrt[t]{C_{1}}}.$
Now, if $y\geq 1/2$ then clearly taking $C_{1}=\log_{3}(5)^{t}$ works. Suppose that instead $y\in(0,1/2)$. Then, since $\log$ grows slower than any polynomial, there exists a $t$-dependent constant $c_{t}<\infty$ such that $\log(3+1/y)\leq c_{t}y^{-1/(2t)}$ for all $y\in(0,1/2)$. Therefore, we have
$3+\frac{1}{y\log(3+1/y)^{t}}\geq 3+\frac{1}{c_{t}^{t}y^{1/2}}.$
It is then clear that
$3+\frac{1}{y}\leq\left(3+\frac{1}{c_{t}^{t}y^{1/2}}\right)^{\sqrt[t]{C_{1}}}$
holds for all $y\in(0,1/2)$ if we take $C_{1}$ large enough in terms of $t$. The second implication follows analogously and we omit its proof. ∎
Lemma 5.
Let $\mu$ be a probability distribution on $\mathbb{R}^{d}$ and $\gamma\in(0,2)$. Then
$\mathbb{E}_{X\sim\mu}\int_{\mathbb{R}^{d}}\frac{(\cos\langle\omega,X\rangle-1)^{2}+\sin^{2}\langle\omega,X\rangle}{\|\omega\|^{d+\gamma}}\mathrm{d}\omega\leq\frac{16\pi^{d/2}M_{\gamma}(\mu)}{\Gamma(d/2)\gamma(2-\gamma)}.$
Proof.
We use the inequalities $(\cos t-1)^{2}+\sin^{2}(t)\leq 4(t^{2}\land 1)$ valid for all $t\in\mathbb{R}$. Plugging in and using the Cauchy-Schwarz inequality, the quantity on the left hand side can be bounded as
$4\mathbb{E}\int_{\mathbb{R}^{d}}\frac{1\land(\|\omega\|^{2}\|X\|^{2})}{\|\omega\|^{d+\gamma}}\mathrm{d}\omega$ $\leq 4\operatorname{vol}_{d-1}(\mathbb{S}^{d-1})\mathbb{E}\int_{0}^{\infty}\frac{1\land(r^{2}\|X\|^{2})}{r^{1+\gamma}}\mathrm{d}r$
$=\frac{8\pi^{d/2}}{\Gamma(\frac{d}{2})}\mathbb{E}\left\{\|X\|^{2}\int_{0}^{\|X\|^{-1}}\frac{1}{r^{\gamma-1}}\mathrm{d}r+\int_{\|X\|^{-1}}^{\infty}\frac{1}{r^{1+\gamma}}\mathrm{d}r\right\}$
$=\frac{16\pi^{d/2}M_{\gamma}(\nu)}{\Gamma(\frac{d}{2})\gamma(2-\gamma)},$
where $\operatorname{vol}_{d-1}(\mathbb{S}^{d-1})={\frac{2\pi^{d/2}}{\Gamma{\bigl{(}}{\frac{d}{2}}{\bigr{)}}}}$ is the surface area of the unit $(d-1)$-sphere. ∎
Lemma 6.
For $\gamma\in(0,2)$ define
$B_{\gamma}=\begin{cases}\sup_{0<a<c}\left|\int_{a}^{c}\frac{\sin(\omega)}{\omega}\mathrm{d}{\omega}\right|&\text{if }\gamma=1,\\ \sup_{0<a<c}\left|\int_{a}^{c}\frac{\cos(\omega)}{\omega^{(1+\gamma)/2}}\mathrm{d}{\omega}\right|&\text{if }\gamma\in(0,1),\\ \sup_{0<a<c}\left|\int_{a}^{c}\frac{\cos(\omega)-1}{\omega^{(1+\gamma)/2}}\mathrm{d}{\omega}\right|&\text{if }\gamma\in(1,2).\end{cases}$
Then $B_{\gamma}<\infty$.
Proof.
In the $\gamma>1$ case, one immediately has $B_{\gamma}\leq\int_{0}^{\infty}\left|\frac{\min\{1,\omega^{2}/2\}}{\omega^{(1+\gamma)/2}}\right|\mathrm{d}{\omega}<\infty.$ Now let us consider $\gamma\leq 1$. One has that $B_{\gamma}=\sup_{0<c_{1}<c_{2}}|I_{\gamma}(c_{2})-I_{\gamma}(c_{1})|\leq\sup_{c>0}2|I_{\gamma}(c)|$ where
$I_{\gamma}(a)=\begin{cases}\int_{1}^{a}\frac{\sin(\omega)}{\omega}\mathrm{d}{\omega}&\text{if }\gamma=1,\\ \int_{1}^{a}\frac{\cos(\omega)}{\omega^{(1+\gamma)/2}}\mathrm{d}{\omega}&\text{if }\gamma\in(0,1),\end{cases}$
Clearly $a\mapsto I_{\gamma}(a)$ is continuous on $(0,\infty)$ for each $\gamma\in(0,2)$. Moreover,
$|I_{\gamma}(a)|\leq|a-1|\cdot\max\{a^{-(\gamma+1)/2},1\}.$
Therefore, we only need to show that $\lim_{a\to\infty}|I_{\gamma}(a)|$ and $\lim_{a\to 0}|I_{\gamma}(a)|$ are both finite. Let $g_{\gamma}(x)=x^{-(\gamma+1)/2}$ and $f_{\gamma}(x)=\sin(x)$ if $\gamma=1$ and $\cos(x)$ otherwise, so that we may write $I_{\gamma}(a)=\int_{a}^{1}f_{\gamma}(\omega)g_{\gamma}(\omega)\mathrm{d}\omega$.
  1. 1.
    For ${a\to\infty}$, since $g_{\gamma}(\infty)=0$ and $|\int_{1}^{a}f_{\gamma}(x)\mathrm{d}t|\leq 2$ is uniformly bounded, $\int_{1}^{\infty}f(\omega)g(\omega)\mathrm{d}\omega$ converges to a finite value by Dirichlet’s test for improper integrals [MA92, page 391] [^6] Reference pointed out by user Siminore on math.stackexchange.com..
  2. 2.
    For $a\to 0$, the conclusion follows by the inequality $|\sin(\omega)|\leq\min\{|\omega|,1\}$ in the case $\gamma=1$, and by $|I_{\gamma}(a)|\leq\int_{a}^{1}\omega^{-(1+\gamma)/2}\mathrm{d}\omega\leq\int_{0}^{1}\omega^{-(1+\gamma)/2}\mathrm{d}\omega=\frac{2}{1-\gamma}$ for $\gamma<1$.
Therefore, $\sup_{a>0}|I_{\gamma}(a)|<\infty$ which concludes the proof.
Lemma 7.
Let $\int_{0}^{\infty}\cdot\,\mathrm{d}\omega\triangleq\lim_{\epsilon\to 0}\int_{1/\epsilon\geq\omega\geq\epsilon}\cdot\,\mathrm{d}\omega$ and recall the definition of $\psi_{\gamma}$ from (2.6). Then, for $x\neq 0$ the following hold:
$\psi_{\gamma}(x)=C_{\psi_{\gamma}}\begin{cases}\int_{0}^{\infty}\frac{\sin(\omega x)}{\omega}\mathrm{d}\omega+\frac{\pi}{2}&\text{for }\gamma=1,\\ \int_{0}^{\infty}\frac{\cos(\omega x)}{\omega^{(1+\gamma)/2}}\mathrm{d}\omega&\text{for }\gamma\in(0,1),\\ \int_{0}^{\infty}\frac{\cos(\omega x)-1}{\omega^{(1+\gamma)/2}}\mathrm{d}\omega&\text{for }\gamma\in(1,2),\end{cases}$
where
$C_{\psi_{\gamma}}=\begin{cases}\left(\cos(\frac{\pi(\gamma-1)}{4})\Gamma(\frac{1-\gamma}{2})\right)^{-1}&\text{if }\gamma\neq 1,\\ \frac{1}{\pi}&\text{if }\gamma=1.\end{cases}$
Proof.
For $x\neq 0$ clearly
$\int_{0}^{\infty}\frac{\sin(\omega x)}{w}\mathrm{d}{w}=\operatorname{sign}(x)\,\int_{0}^{\infty}\frac{\sin(\omega)}{w}\mathrm{d}{w}=\operatorname{sign}(x)\frac{\pi}{2},$
which shows the first claim. Assume from here on without loss of generality that $x>0$. For $\gamma\in(0,1)$, by the residue theorem,
$\int_{0}^{\infty}\frac{\cos(\omega x)}{\omega^{(1+\gamma)/2}}\mathrm{d}{\omega}$ $=x^{(\gamma-1)/2}\,\int_{0}^{\infty}\Re\left(\frac{e^{i\omega}}{\omega^{(1+\gamma)/2}}\right)\mathrm{d}{w}$
$=x^{(\gamma-1)/2}\,\Re\left(ie^{-i\frac{\pi}{2}\gamma}\right)\int_{0}^{\infty}\frac{e^{-z}}{z^{(1+\gamma)/2}}\mathrm{d}{z}$
$=x^{(\gamma-1)/2}\,\cos\left(\frac{\pi(\gamma-1)}{4}\right)\Gamma((1-\gamma)/2).$
Similarly, for $\gamma\in(1,2)$, integration by parts and the residue theorem gives
$\int_{0}^{\infty}\frac{\cos(\omega x)-1}{\omega^{(1+\gamma)/2}}\mathrm{d}{\omega}$ $=x^{(\gamma-1)/2}\,\int_{0}^{\infty}(\cos(\omega)-1)\mathrm{d}\left(\frac{-1}{((\gamma-1)/2)\,\omega^{(\gamma-1)/2}}\right)$
$=-x^{(\gamma-1)/2}\int_{0}^{\infty}\frac{\sin(w)}{(\gamma-1)/2\,\omega^{(\gamma-1)/2}}\mathrm{d}{\omega}$
$=-x^{(\gamma-1)/2}\frac{2}{\gamma-1}\int_{0}^{\infty}\mathfrak{I}\left(\frac{e^{iw}}{\omega^{(\gamma-1)/2}}\right)\mathrm{d}{\omega}$
$=-x^{(\gamma-1)/2}\frac{2}{\gamma-1}\mathfrak{I}\left(ie^{-\frac{\pi}{2}(\gamma-1)/2}\right)\int_{0}^{\infty}\frac{e^{-z}}{z^{(\gamma-1)/2}}\mathrm{d}{z}$
$=-x^{(\gamma-1)/2}\frac{2}{\gamma-1}\cos\left(\frac{\pi(\gamma-1)}{4}\right)\Gamma(1-(\gamma-1)/2)$
$=x^{(\gamma-1)/2}\,\cos\left(\frac{\pi(\gamma-1)}{4}\right)\Gamma((1-\gamma)/2).$
Lemma 8.
Let $\phi$ be the probability density function of $\mathcal{N}(0,\sigma I_{d})$ and write $\widehat{\phi}$ for its Fourier transform. Then, for any $\beta\geq 1$,
$\|\widehat{\phi}(\omega)\|\omega\|^{\beta}\|_{2}^{2}=\frac{\pi^{d/2}}{\Gamma(d/2)\sigma^{2\beta+d}}\Gamma\left(\frac{2\beta+d}{2}\right)\leq\frac{5\pi^{d/2}}{\Gamma(d/2)\sigma^{2\beta+d}}\left(\frac{2\beta+d}{2e}\right)^{\frac{2\beta+d-1}{2}}.$
Proof.
It is well known that $\widehat{\phi}(\omega)=e^{-\frac{\sigma^{2}}{2}\|\omega\|^{2}}$. The claimed equality then follows from the formula for the $2\beta$’th moment of the Gaussian distribution with mean 00 and variance $1/(2\sigma^{2})$. The inequality follows by Item 2 of Lemma 9.
Lemma 9 (Properties of the gamma function).
The $\Gamma$ function has the following properties.
  1. 1.
    For all $x>0$ and $s\in(0,1)$ we have $\Gamma(x+s)\leq x^{s}\Gamma(x)$.
  2. 2.
    For all $x>1$ we have $\Gamma(x)\leq 5(x/e)^{x-1/2}$.
  3. 3.
    For all $x>0$ we have $x\Gamma(x)\geq 0.885$.
Proof.
For the first claim see [Wen48] [^7] Reference pointed out by Egor Maximenko on math.stackexchange.com. In [MS64] authors showed that $\log\Gamma(x)\leq(x-\frac{1}{2})\log(x)-x+\frac{1}{2}\log(2\pi)+1$ for all $x\geq 1$, from which the second claim follows as $\exp(\frac{1}{2}\log(2\pi)+1/2)<5$. The third claim can be verified numerically. ∎
Lemma 10.
Let $b\geq 1$ and $|a|<b$. Then
$\int_{0}^{r}x^{a}|\sin(x)|^{b}\mathrm{d}x=O(r^{a+b})$
as $r\to\infty$, where we hide constants depending on $a,b$.
Proof.
Since we are only interested in the asymptotic behaviour as $r\to\infty$, assume without loss of generality that $r\geq 1$. Then, we have
$\int_{0}^{r}x^{a}|\sin(x)|^{b}\mathrm{d}x=\underbrace{\int_{0}^{1}x^{a}|\sin(x)|^{b}\mathrm{d}x}_{I}+\underbrace{\int_{1}^{r}x^{a}|\sin(x)|^{b}\mathrm{d}x}_{II}.$
Using the inequality $|\sin(x)|\leq x$, we can bound the first term as
$I\leq\int_{0}^{1}x^{a+b}\mathrm{d}x=\frac{1}{a+b+1}\leq 1=O(r^{a+b}),$
since $a+b>0$. For the second term, we obtain
$II\leq\int_{1}^{r}x^{a}\mathrm{d}x=\begin{cases}\begin{rcases}\frac{r^{a+1}-1}{a+1}&\text{if }a\neq-1\\ \log(r)&\text{if }a=-1\end{rcases}\end{cases}=O(r^{a+b}),$
where the last step uses $a+b>0$ and $b\geq 1$.
Lemma 11.
Let $a,b,c\in\mathbb{R}$ with $b>0$ be constants. For all large enough $r$ one has
$\int_{r}^{\infty}x^{a}\exp\left(-\frac{bx}{\log^{2}(x+2)}\right)\mathrm{d}x<r^{-c}.$
Proof.
Assume, without loss of generality, that $c\geq 0$. For all large enough $x$ one has
$\exp\left(-\frac{bx}{\log^{2}(x+2)}\right)<x^{-a-c-2}.$
Therefore, for large enough $r$,
$\int_{r}^{\infty}x^{a}\exp\left(-\frac{bx}{\log^{2}(x+2)}\right)\mathrm{d}x<\int_{r}^{\infty}x^{-c-2}\mathrm{d}x\asymp r^{-c-1}<r^{-c}.$
Lemma 12.
Let $J_{\nu}$ be the Bessel function of the first kind of order $\nu$.
  1. 1.
    For all $x\in\mathbb{R}^{d}$,
    $\int_{\mathbb{R}^{d}}e^{i\langle x,v\rangle}\mathrm{d}\sigma(v)=(2\pi)^{d/2}\|x\|^{1-d/2}J_{d/2-1}(\|x\|).$
  2. 2.
    For any $\nu\in\mathbb{R}$, as $x\to\infty$
    $J_{\nu}(x)=\sqrt{\frac{2}{\pi x}}\cos\left(x-{\frac{\nu\pi}{2}}-\frac{\pi}{4}\right)+O(x^{-3/2}).$ (A.1)
  3. 3.
    For all $x\in\mathbb{R}^{d}$,
    $\int_{\mathbb{B}^{d}(0,1)}e^{i\langle x,w\rangle}\mathrm{d}w=\left(\frac{2\pi}{\|x\|}\right)^{d/2}J_{d/2}(\|x\|).$
Proof.
For Item 1 set $r=\|x\|$ and $s=(2\pi)^{-1}$ in the calculation [SW71, page 154].
For Item 2 see [Wat80, Equation (1) in Section 7.21].
For Item 3, we can compute
$\int_{\|w\|\leq 1}e^{i\langle x,w\rangle}\mathrm{d}w$ $=\int_{-1}^{1}e^{i\|x\|w_{1}}\int_{w_{2}^{2}+\dots+w_{d}^{2}\leq 1-w_{1}^{2}}\mathrm{d}w_{2}\dots\mathrm{d}w_{d}\mathrm{d}w_{1}$
$=\frac{\pi^{(d-1)/2}}{\Gamma(\frac{d+1}{2})}\int_{-1}^{1}e^{i\|x\|w_{1}}(1-w_{1}^{2})^{(d-1)/2}\mathrm{d}w_{1}.$ (A.2)
Recall the definition (from e.g. [Wat80, Section 3.1]) of the Bessel function of the first kind as
$J_{\nu}(x)=\frac{(x/2)^{\nu}}{\Gamma(\nu+\frac{1}{2})\Gamma(\frac{1}{2})}\int_{0}^{\infty}\cos(x\cos(\theta))\sin^{2\nu}(\theta)\mathrm{d}\theta$
valid for $\nu>-1/2$; the above is also known as the Poisson representation. Changing variables to $u=\cos(\theta)$, we see that it is equal to
$J_{\nu}(x)$ $=\frac{(x/2)^{\nu}}{\Gamma(\nu+\frac{1}{2})\Gamma(\frac{1}{2})}\int_{-1}^{1}e^{ixu}(1-u^{2})^{\nu-1/2}\mathrm{d}u.$ (A.3)
Comparing (A.3) with (A) concludes the proof. ∎
Lemma 13.
There exists a radial function $h_{0}\in L^{2}(\mathbb{R}^{d})$ such that
$\operatorname{supp}(h_{0})\subseteq\mathbb{B}(0,1),$
$|\widehat{h}_{0}(w)|\leq C\exp\left(-\frac{c\|w\|}{\log(\|w\|+2)^{2}}\right)$ for all $\quad\text{for all }w\in\mathbb{R}^{d},$
$|\widehat{h}_{0}(w)|\geq\frac{1}{2}$ for all $\quad\text{for all }\|w\|\leq r_{\min},$
where $C,c,r_{\min}>0$.
Proof.
Apply Theorem 1.4 in [Coh23] using the spherically symmetric weight function $u:\mathbb{R}^{d}\to\mathbb{R}_{\leq 0}$ defined by
$u(w)=u(\|w\|)=-\frac{\|w\|}{\log(\|w\|+2)^{2}}\left(\frac{(\|w\|-2)_{+}}{\|w\|+2}\right)^{4},$
where $(a)_{+}:=\max(a,0)$ for $a\in\mathbb{R}$. ∎
Appendix B Proof of Proposition 4
For $v\in\mathbb{S}^{d-1}$ and $b\in\mathbb{R}$ let $\theta_{v}(x)=\langle v,x\rangle$ and write $\eta_{v}\triangleq\theta_{v}\#(\mu-\nu)$ for the pushforward of the measure $\mu-\nu$ through the map $\theta_{v}$. To start with, we notice that
$\int_{\mathbb{S}^{d-1}}\int_{\mathbb{R}}\Big{[}\mathbb{E}\psi_{\gamma}(\langle X,v\rangle-b)-\mathbb{E}\psi_{\gamma}(\langle Y,v\rangle-b)\Big{]}^{2}\mathrm{d}b\mathrm{d}\sigma(v)$ (B.1)
$=$ $\int_{\mathbb{S}^{d-1}}\int_{\mathbb{R}}\left(\int_{\mathbb{R}}\psi_{\gamma}(x-b)\mathrm{d}\eta_{v}(x)\right)^{2}\mathrm{d}b\mathrm{d}\sigma(v),$
For each $v\in\mathbb{S}^{d-1}$, the measure $\eta_{v}$ has at most countably many atoms, therefore $b\mapsto\eta_{v}(\{b\})=0$ $\operatorname{Leb}$-almost everywhere. Then, by Tonelli’s theorem we can conclude that $\eta_{v}(\{b\})=0$ for $\sigma\otimes\operatorname{Leb}$-almost every $(v,b)$, thus going forward we can focus on the case $x\neq b$. By Lemma 7, and writing $A_{\epsilon}=[\epsilon,1/\epsilon]$ for $\epsilon>0$, we have
$\int_{\mathbb{R}}\psi_{\gamma}(x-b)\mathrm{d}\eta_{v}(x)=C_{\psi_{\gamma}}\int_{\mathbb{R}}\lim_{\varepsilon\to 0}\int_{A_{\varepsilon}}\begin{rcases}\begin{dcases}\frac{\sin(\omega(x-b))}{\omega}&\text{if }\gamma=1\\ \frac{\cos(\omega(x-b))}{\omega^{(1+\gamma)/2}}&\text{if }\gamma\in(0,1)\\ \frac{\cos(\omega(x-b))-1}{\omega^{(1+\gamma)/2}}&\text{if }\gamma\in(1,2)\end{dcases}\end{rcases}\mathrm{d}\omega\mathrm{d}\eta_{v}(x).$
Note that in the $\gamma=1$ case we implicitly used that $\int\mathrm{d}\eta_{v}(x)=0$. To exchange the integral over $x$ and the limit over $\varepsilon$, notice that for any $\varepsilon>0$ and $x\neq b\in\mathbb{R}$,
$\left|\int_{\varepsilon}^{1/\varepsilon}\frac{\sin(\omega(x-b))}{\omega}\mathrm{d}\omega\right|\leq B_{\gamma}$ if $\text{if }\gamma=1,$
$\left|\int_{\varepsilon}^{1/\varepsilon}\frac{\cos(\omega(x-b))}{\omega^{(1+\gamma)/2}}\mathrm{d}\omega\right|\leq B_{\gamma}|x-b|^{(\gamma-1)/2}$ if $\text{if }\gamma\in(0,1),$
$\left|\int_{\varepsilon}^{1/\varepsilon}\frac{\cos(\omega(x-b))-1}{\omega^{(1+\gamma)/2}}\mathrm{d}\omega\right|\leq B_{\gamma}|x-b|^{(\gamma-1)/2}$ if $\text{if }\gamma\in(1,2).$
where $B_{\gamma}<\infty$ depends only on $\gamma$ and is defined in Lemma 6. We now show that $\int_{\mathbb{R}}|x-b|^{(\gamma-1)/2}\mathrm{d}|\eta_{v}|(x)<\infty$ for $\sigma\otimes\operatorname{Leb}$-almost every $b,v$. To this end, let $S=\{(b,v)\in\mathbb{R}\times\mathbb{S}^{d-1}:\int_{\mathbb{R}}|x-b|^{(\gamma-1)/2}\mathrm{d}|\eta_{v}|(x)=\infty\}$ and assume for contradiction $(\sigma\otimes\operatorname{Leb})(S)>0$. Then $\mathbbm{1}_{([-B,B]\times\mathbb{S}^{d-1})\cap S}\uparrow\mathbbm{1}_{S}$ as $B\to\infty$, and thus by the monotone convergence theorem there exists a finite $B$ such that $\operatorname{Leb}(([-B,B]\times\mathbb{S}^{d-1})\cap S)>0$. However, by Tonelli’s theorem we have
$\int_{-B}^{B}\left(\int_{\mathbb{R}}|x-b|^{(\gamma-1)/2}\mathrm{d}|\eta_{v}|(x)\right)^{2}\mathrm{d}b$ $\leq\int_{-B}^{B}\int_{\mathbb{R}}|x-b|^{\gamma-1}\mathrm{d}|\eta_{v}|(x)\mathrm{d}b$
$\leq 2\int_{\mathbb{R}}\int_{0}^{B+|x|}b^{\gamma-1}\mathrm{d}b\mathrm{d}|\eta_{v}|(x)$
$\lesssim\int_{\mathbb{R}}(B+|x|)^{\gamma}\mathrm{d}|\eta_{v}|(x)$
$\lesssim B^{\gamma}+\mathbb{E}_{X\sim\mu}\left[|\langle v,X\rangle|^{\gamma}\right]+\mathbb{E}_{Y\sim\nu}\left[|\langle v,Y\rangle|^{\gamma}\right]$
$\leq B^{\gamma}+M_{\gamma}(\mu+\nu),$
which, after integration over $v\in\mathbb{S}^{d-1}$, leads to a contradiction if $M_{\gamma}(\mu+\nu)<\infty$. Continuing under the assumption $M_{\gamma}(\mu+\nu)<\infty$, we can apply the dominated convergence theorem to obtain
$\int_{\mathbb{R}}\psi_{\gamma}(x-b)\mathrm{d}\eta_{v}(x)=C_{\psi_{\gamma}}\lim_{\varepsilon\to 0}\int_{\mathbb{R}}\int_{A_{\varepsilon}}\begin{rcases}\begin{dcases}\frac{\sin(\omega(x-b))}{\omega}&\text{if }\gamma=1\\ \frac{\cos(\omega(x-b))}{\omega^{(1+\gamma)/2}}&\text{if }\gamma\in(0,1)\\ \frac{\cos(\omega(x-b))-1}{\omega^{(1+\gamma)/2}}&\text{if }\gamma\in(1,2)\end{dcases}\end{rcases}\mathrm{d}\omega\mathrm{d}\eta_{v}(x).$
Then by Fubini’s theorem, we exchange the order of integration to get
$\int_{\mathbb{R}}\psi_{\gamma}(x-b)\mathrm{d}\eta_{v}(x)=C_{\psi_{\gamma}}\lim_{\varepsilon\to 0}\int_{A_{\varepsilon}}\int_{\mathbb{R}}\begin{rcases}\begin{dcases}\frac{\sin(\omega(x-b))}{\omega}&\text{if }\gamma=1\\ \frac{\cos(\omega(x-b))}{\omega^{(1+\gamma)/2}}&\text{if }\gamma\in(0,1)\\ \frac{\cos(\omega(x-b))-1}{\omega^{(1+\gamma)/2}}&\text{if }\gamma\in(1,2)\end{dcases}\end{rcases}\mathrm{d}\eta_{v}(x)\mathrm{d}\omega.$
Notice that $\int_{\mathbb{R}}e^{-i\omega x}\mathrm{d}\eta_{v}(x)=\widehat{\eta}_{v}(\omega)$, $\widehat{\eta}_{v}(\omega)=\overline{\widehat{\eta}_{v}(-\omega)}$ and $\widehat{\eta}_{v}(0)=0$,
$\int_{\mathbb{R}}\psi_{\gamma}(x-b)\mathrm{d}\eta_{v}(x)$ $=C_{\psi_{\gamma}}\lim_{\varepsilon\to 0}\int_{A_{\varepsilon}}\frac{1}{\omega^{(1+\gamma)/2}}\begin{rcases}\begin{dcases}\mathfrak{I}(e^{-i\omega b}\overline{\widehat{\eta}_{v}(\omega)})&\text{if }\gamma=1\\ \Re(e^{-i\omega b}\overline{\widehat{\eta}_{v}(\omega)})&\text{if }\gamma\neq 1\end{dcases}\end{rcases}\mathrm{d}\omega$
$=C_{\psi_{\gamma}}\lim_{\varepsilon\to 0}\begin{cases}\mathfrak{I}\left(\widehat{\Psi}_{\gamma,v,\varepsilon}(b)\right)&\text{if }\gamma=1\\ \Re\left(\widehat{\Psi}_{\gamma,v,\varepsilon}(b)\right)&\text{if }\gamma\neq 1.\end{cases}$
where we write
$\Psi_{\gamma,v,\varepsilon}(\omega)=\frac{\overline{\widehat{\eta}_{v}(\omega)}}{\omega^{(1+\gamma)/2}}\mathbbm{1}\{\omega\in A_{\varepsilon}\}.$
Notice that $\Psi_{\gamma,v,\varepsilon}$ is bounded and compactly supported and thus lies in $L^{p}(\mathbb{R})$ for any $p$, and so in particular
$\Psi_{\gamma,v,\varepsilon}\in L^{1}(\mathbb{R})\cap L^{2}(\mathbb{R}),$
which ensures that
$\widehat{\Psi}_{\gamma,v,\varepsilon}\in L^{\infty}(\mathbb{R})\cap L^{2}(\mathbb{R}).$
Finally, let us write
$\Psi_{\gamma,v}(\omega)=\lim_{\epsilon\to 0}\Psi_{\gamma,v,\varepsilon}(\omega)=\frac{\overline{\widehat{\eta}_{v}(\omega)}}{\omega^{(1+\gamma)/2}}\mathbbm{1}\{\omega>0\}$
for every $\omega$. We now show that $\Psi_{\gamma,v}\in L^{2}(\mathbb{R})$ provided $M_{\gamma}(\mu+\nu)<\infty$, which is assumed throughout. Let $(X,Y)\sim\mu\otimes\nu$. We have
$\int_{\mathbb{R}}|\Psi_{\gamma,v}(\omega)|^{2}\mathrm{d}\omega$ $=\int_{0}^{\infty}\frac{|\widehat{\eta}_{v}(\omega)|^{2}}{w^{1+\gamma}}\mathrm{d}w$
$=\int_{0}^{\infty}\frac{(\mathbb{E}[\cos\langle\omega,X\rangle-\cos\langle\omega,Y\rangle])^{2}+(\mathbb{E}[\sin\langle\omega,X\rangle-\sin\langle\omega,Y\rangle])^{2}}{\omega^{1+\gamma}}\mathrm{d}\omega.$
Using the inequality $(a-b)^{2}\leq 2(a-1)^{2}+(b-1)^{2},\,\forall\,a,b\in\mathbb{R}$ for the $\cos$ term, the inequality $(a+b)\leq 2a^{2}+2b^{2},\,\forall\,a,b\in\mathbb{R}$ for the $\sin$ term, and applying Jensen’s inequality to take the expectation outside, we can conclude that $\Psi_{\gamma,v}\in L^{2}(\mathbb{R})$ by Lemma 5. Thus, by the dominated convergence theorem
$\left\|\Psi_{\gamma,v,\epsilon}-\Psi_{\gamma,v}\right\|_{2}\to 0$
as $\epsilon\to 0$. Then, by Parseval’s identity
$\left\|\widehat{\Psi}_{\gamma,v,\epsilon}-\widehat{\Psi}_{\gamma,v}\right\|_{2}\to 0$ (B.2)
as $\epsilon\to 0$. It is well know that convergence in $L^{2}(\mathbb{R})$ implies that there exists a subsequence $\{\varepsilon_{n}\}_{n=1}^{\infty}$ with $\varepsilon_{n}\to 0$ and $\widehat{\Psi}_{\gamma,v,\epsilon}\to\widehat{\Psi}_{\gamma,v}$ almost everywhere888we could also conclude this by Carleson’s theorem. Therefore, by passing to this subsequence, it follows that
$\int_{\mathbb{R}}\psi_{\gamma}(x-b)\mathrm{d}\eta_{v}(x)=C_{\psi_{\gamma}}\begin{cases}\mathfrak{I}\left(\widehat{\Psi}_{\gamma,v}(b)\right)&\text{if }\gamma=1\\ \Re\left(\widehat{\Psi}_{\gamma,v}(b)\right)&\text{if }\gamma\neq 1\end{cases}$
for $\sigma\otimes\operatorname{Leb}$-almost every $(b,v)\in\mathbb{R}\times\mathbb{S}^{d-1}$. Note that since $\eta_{v}(\omega)\in\mathbb{R}$,
$\Re\left(\widehat{\Psi}_{\gamma,v}(b)\right)$ $=\frac{\widehat{\Psi}_{\gamma,v}(b)+\overline{\widehat{\Psi}_{\gamma,v}(b)}}{2}$
$=\frac{1}{2}\int_{0}^{\infty}\left(\frac{\widehat{\eta}_{v}(\omega)}{\omega^{(1+\gamma)/2}}e^{ib\omega}+\frac{\widehat{\eta}_{v}(-\omega)}{\omega^{(1+\gamma)/2}}e^{-ib\omega}\right)\mathrm{d}\omega$
$=\frac{1}{2}\int_{-\infty}^{\infty}\frac{\widehat{\eta}_{v}(\omega)\operatorname{sign}(\omega)}{|\omega|^{(1+\gamma)/2}}e^{ib\omega}\mathrm{d}\omega$
$=\mathcal{F}\left[\frac{\widehat{\eta}_{v}(\omega)\operatorname{sign}(\omega)}{2|\omega|^{(1+\gamma)/2}}\right](-b),$ (B.3)
$\mathfrak{I}\left(\widehat{\Psi}_{\gamma,v}(b)\right)$ $=\frac{\widehat{\Psi}_{\gamma,v}(b)-\overline{\widehat{\Psi}_{\gamma,v}(b)}}{2i}$
$=\frac{1}{2i}\int_{0}^{\infty}\left(\frac{\overline{\widehat{\eta}_{v}(\omega)}}{\omega^{(1+\gamma)/2}}e^{-ib\omega}-\frac{\overline{\widehat{\eta}_{v}(-\omega)}}{\omega^{(1+\gamma)/2}}e^{ib\omega}\right)\mathrm{d}\omega$
$=\frac{1}{2i}\int_{-\infty}^{\infty}\frac{\overline{\widehat{\eta}_{v}(\omega)}}{|\omega|^{(1+\gamma)/2}}\operatorname{sign}(\omega)e^{ib\omega}\mathrm{d}\omega$
$=\mathcal{F}\left[\frac{\overline{\widehat{\eta}_{v}(\omega)}\operatorname{sign}(\omega)}{2i|\omega|^{(1+\gamma)/2}}\right](-b).$ (B.4)
Plugging (B.3) and (B.4) into (B.1), by Parseval’s identity (implicitly using that $\Psi_{\gamma,v}\in L^{2}(\mathbb{R})$), we obtain
$\int_{\mathbb{S}^{d-1}}\int_{\mathbb{R}}\Big{[}\mathbb{E}\psi_{\gamma}(\langle X,v\rangle-b)-\mathbb{E}\psi_{\gamma}(\langle Y,v\rangle-b)\Big{]}^{2}\mathrm{d}b\mathrm{d}\sigma(v)$ $=\int_{\mathbb{S}^{d-1}}\int_{\mathbb{R}}\left(\int_{\mathbb{R}}\psi_{\gamma}(x-b)\mathrm{d}\eta_{v}(x)\right)^{2}\mathrm{d}b\mathrm{d}\sigma(v),$
$=2\pi C_{\psi_{\gamma}}^{2}\int_{\mathbb{S}^{d-1}}\int_{\mathbb{R}}\frac{|\widehat{\eta}_{v}(\omega)|^{2}}{4|\omega|^{1+\gamma}}\mathrm{d}w\mathrm{d}\sigma(v)$
$=\pi C_{\psi_{\gamma}}^{2}\int_{\mathbb{S}^{d-1}}\int_{0}^{\infty}\frac{|\widehat{\eta}_{v}(\omega)|^{2}}{|\omega|^{1+\gamma}}\mathrm{d}w\mathrm{d}\sigma(v)$
$=\pi C_{\psi_{\gamma}}^{2}\int_{\mathbb{R}^{d}}\frac{|\mathcal{F}[\mu-\mathcal{\nu}](\omega)|^{2}}{\|\omega\|^{d+\gamma}}\mathrm{d}\omega,$
where the last step uses a polar change of variable. The result follows after comparing with Proposition 3.
Appendix C Proof of Proposition 5
Let $\mathcal{S}(\mathbb{R}^{d})$ be the Schwartz space and $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ be the space of all tempered distributions on $\mathbb{R}^{d}$. Let $\tau=\mu-\nu$ and $s=(d+\gamma)/2$. First, note that
$\int\frac{K_{s}(x)\mathrm{d}{x}}{(1+\|x\|^{2})^{d}}<\infty,$
so by [Lan72, Theorem 0.10] we have $K_{s}\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$. By [Lan72, Theorem 0.12], since $K_{s}\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$ and $\tau$ has compact support,
$\widehat{I_{s}f}=\widehat{K_{s}*\tau}=\widehat{K_{s}}\widehat{\tau}.$
By Plancherel’s identity,
$(2\pi)^{\frac{d}{2}}\|I_{s}\tau\|_{2}=\|\widehat{I_{s}\tau}\|_{2}=\|\widehat{K_{s}}\widehat{\tau}\|_{2}=\frac{1}{\sqrt{F_{\gamma}(d)}}\mathcal{E}_{\gamma}(\mu,\nu),$
where the last equality follows from Proposition 3.
Appendix D Proof of Theorem 7 and Proposition 8
In this section we prove both Proposition 7 and Proposition 8. To do so, we give two constructions. The first one, presented in Section D.1, only applies in one dimension and gives optimal results. The second construction is given in Section D.2 applies in all dimensions, but loses a polylogarithmic factor.
Notation: Abusing notation, in what follows we write $\mathcal{E}_{\gamma}(f,g)$ and $\overline{d_{H}}(f,g)$ even when $f$ and $g$ are not necessarily probability measures or probability densities. We will also write $\|f\|_{t,2}=\|\|\cdot\|^{t}\widehat{f}\|_{2}$ for potentially negative exponents $t\in\mathbb{R}$. Note that $\mathcal{E}_{\gamma}(f,0)=\sqrt{F_{\gamma}(d)}\|\widehat{f}\|_{-\frac{d+\gamma}{2},2}$.
D.1 The case $d=1$
The Lemma below constructs the difference of two densities that has favorable properties.
Lemma 14.
Let $f(x)=1\{|x|\leq\pi\}\sin(rx)$ with $r\in\mathbb{Z}$ and write $f_{\beta}=f*\cdots*f$ for $f$ convolved with itself $\beta-1$ times, i.e. $f_{1}=f,f_{2}=f*f$ and so on. Fix an integer $\beta\geq 1$ and let $|t|<\beta$. We have
$\|f_{\beta}\|_{t,2}\asymp r^{t},\|f_{\beta}\|_{1}\asymp 1,\text{ and }\overline{d_{H}}(f_{\beta},0)\asymp\frac{1}{r},$ (D.1)
as $r\to\infty$ where the constants may depend on $\beta,t$.
Proof.
The intuition for the estimates (D.1) is simple: most of the energy of $f$ (and hence $f_{\beta}$) is at frequencies around $|\omega|\approx r$ and thus differentiating $t$ times boosts the $L_{2}$-energy by $r^{t}$. A simple computation shows $\widehat{f}(\omega)=c{(-1)^{r}\over i}{r\over\omega^{2}-r^{2}}\sin(\omega\pi)$. Note that because $r\in\mathbb{Z}$ we have $\|\widehat{f}\|\asymp\|\widehat{f}_{\beta}\|\asymp 1$.
Estimating $\|f_{\beta}\|_{t,2}$. By definition we have
$\|f_{\beta}\|^{2}_{t,2}\asymp\int_{0}^{\infty}|\widehat{f}(\omega)|^{2\beta}\omega^{2t}\,\mathrm{d}\omega\asymp\int_{0}^{\infty}{r^{2\beta}\over(\omega^{2}-r^{2})^{2\beta}}\omega^{2t}\sin^{2\beta}(\omega\pi)\mathrm{d}\omega.$
We decompose the integral into three regimes:
  1. 1.
    $\omega<r/2$: here $(\omega^{2}-r^{2})\asymp r^{2}$ and thus
    $\int_{0}^{r/2}(\dots)\asymp r^{-2\beta}\int_{0}^{r/2}\omega^{2t}\sin^{2\beta}(\omega\pi)\lesssim r^{2t}$
    by Lemma 10.
  2. 2.
    $\omega>3r/2$: here $(\omega^{2}-r^{2})\asymp\omega^{2}$ and thus
    $\int_{3r/2}^{\infty}(\dots)\asymp r^{2\beta}\int_{3r/2}^{\infty}{\sin^{2\beta}(\omega\pi)\omega^{2t}\over\omega^{4\beta}}\asymp r^{2\beta}r^{2t-4\beta+1}=r^{1+2t-2\beta}\ll r^{2t}.$
  3. 3.
    $\omega\in[r/2,3r/2]$: here $(\omega^{2}-r^{2})\asymp yr$, where $y=\omega-r$. Note also $\sin(\omega\pi)=\sin(r\pi+y\pi)=(-1)^{r}\sin(y\pi)$, and $\omega\asymp r$. Thus
    $\int_{r/2}^{3r/2}(\dots)\mathrm{d}\omega=\int_{-r/2}^{r/2}(\dots)\mathrm{d}y\asymp r^{2\beta}\int_{-r/2}^{r/2}{\sin^{2\beta}(y\pi)r^{2t}\over(yr)^{2\beta}}\mathrm{d}y\asymp r^{2t}\int_{\mathbb{R}}\left(\frac{\sin(y\pi)}{y}\right)^{2\beta}\mathrm{d}y\asymp r^{2t}.$
    where the last inequality follows by that the integrand is bounded at 00 and has $y^{-2\beta}\lesssim y^{-2}$ tail.
Estimating $\|f_{\beta}\|_{1}$. Follows from $\|f_{\beta}\|_{1}\lesssim\|f_{\beta}\|_{2}\asymp 1$ by the Cauchy-Schwartz inequality and $\|f_{\beta}\|_{1}\geq\|\widehat{f_{\beta}}\|_{\infty}\asymp 1$ by the Hausdorff–Young inequality.
Estimating $\overline{d_{H}}$. We get $\overline{d_{H}}(f_{\beta},0)\gtrsim\mathcal{E}_{1}(f_{\beta},0)\asymp\|f_{\beta}\|_{-1,2}\asymp\frac{1}{r}$ from the first estimate. For the upper bound, note that $\widehat{\operatorname{sign}(x)}=\frac{2}{i\omega}$ and $\overline{d_{H}}(f_{\beta},0)=\sup_{b}\frac{1}{2}\int f_{\beta}(x)\operatorname{sign}(x-b)\mathrm{d}{x}$, so by Plancherel’s identity,
$\overline{d_{H}}(f_{\beta},0)\lesssim\sup_{b}\int\left|\widehat{f_{\beta}}(\omega)\frac{e^{ib\omega}}{\omega}\right|\mathrm{d}{\omega}\lesssim\int_{0}^{\infty}{r^{\beta}\over(\omega^{2}-r^{2})^{\beta}}\omega^{-1}\sin^{\beta}(\omega\pi).$
The fact that the above is $O(1/r)$ follows analogously to the proof of our bound on $\|f_{\beta}\|_{t,2}$ so we omit it. This concludes our proof.
Proof of Proposition 7 and Proposition 8 for $d=1$.
We now turn to showing tightness in one dimension, utilizing the density difference constructed in Lemma 14. Given a value of the smoothness $\beta>0$, set $\overline{\beta}=\lceil\beta\rceil+1$ and let $f_{\overline{\beta}}$ be as in Lemma 14 with $r=\epsilon^{-1/\beta}$ for some $\epsilon\in(0,1)$. Let $p_{0}$ be a smooth, compactly supported density with $\inf_{x\in[-\pi,\pi]}p_{0}(x)>0$. Define
$p_{\epsilon}(x)=p_{0}(x)+\epsilon f_{\overline{\beta}}(\overline{\beta}x)/2\qquad\text{and}\qquad q_{\epsilon}(x)=p_{0}(x)-\epsilon f_{\overline{\beta}}(\overline{\beta}x)/2.$
Clearly both $p_{\epsilon},q_{\epsilon}$ are compactly supported probability densities for sufficiently small $\epsilon$, since $\|f_{\overline{\beta}}\|_{\infty}<\infty$ and is supported on $[-\overline{\beta}\pi,\overline{\beta}\pi]$. By Lemma 14, for each $\gamma\in(0,2)$ the two densities satisfy
$\|p_{\epsilon}-q_{\epsilon}\|_{1}\asymp\epsilon,\,\|p_{\epsilon}\|_{\beta,2}\asymp\|q_{\epsilon}\|_{\beta,2}\asymp 1,\,\mathcal{E}_{\gamma}(p_{\epsilon},q_{\epsilon})\asymp\|p_{\epsilon}-q_{\epsilon}\|_{-(1+\gamma)/2,2}\asymp\epsilon^{\frac{2\beta+\gamma+1}{2\beta}},\overline{d_{H}}(p_{\epsilon},q_{\epsilon})\asymp\epsilon^{\frac{\beta+1}{\beta}}.$
This proves both Proposition 7 and Proposition 8 for $d=1$. ∎
D.2 The case $d>1$
D.2.1 Overview
For the discussions below, we will assume that the ambient dimension $d\geq 2$. Our construction here is less straightforward than for $d=1$ in Section D.1 but shares the same basic idea. Recall that the basic premise is that we want to saturate the Hölder’s inequality in 3.3, which requires the density difference $f=\mu-\nu$ to have Fourier transform be (almost) supported on a sphere. For $d=1$ we took $f$ to be a pure sinusoid. However, of course such $f$ is not compactly supported and that is why we multiplied the sinusoid by a rectangle (and then convolved many times to gain smoothness), which served as a mollifier.
For $d>1$ let us attempt to follow the same strategy and take
$f_{r}(x)=g_{r}(x)h(x)\,,$
where $r>0$ is a parameter, $h$ is some compactly supported smooth mollifier and $g_{r}(x)$ is defined implicitly via
$\widehat{g}_{r}(\omega)=r^{(1-d)/2}\delta(\|\omega\|-r),$
where here and below we denote, a bit informally, by $\delta(\|\cdot\|-r)$ a distribution that integrates any smooth compactly supported function $\phi$ as follows:
$\int_{\mathbb{R}^{d}}\phi(\omega)\delta(\|\omega\|-r)\mathrm{d}\omega\triangleq r^{d-1}\int_{\mathbb{R}^{d}}\phi(r\omega)\mathrm{d}\sigma(\omega)=\frac{2\pi^{d/2}r^{d-1}}{\Gamma(\frac{d}{2})}\mathbb{E}[\phi(rX)]\,,$
where $\sigma$ is the unnormalized surface measure of $\mathbb{S}^{d-1}$ and $X$ is a random vector uniformly distributed on $\mathbb{S}^{d-1}$. Explicit computation shows
$g_{r}(x)$ $=\mathcal{F}^{-1}[\widehat{g}_{r}](x)=\frac{\sqrt{r}}{(2\pi)^{d}r^{d/2}}\int_{\mathbb{R}^{d}}e^{i\langle\omega,x\rangle}\delta(\|\omega\|-r)\mathrm{d}\omega$
$=\frac{\sqrt{r}}{(2\pi)^{d/2}}\,\|x\|^{1-d/2}\,J_{d/2-1}(\|rx\|),$
where $J_{\nu}$ denotes Bessel functions of the first kind of order $\nu$. Notice that $g$ is spherically symmetric and real-valued (some further properties of it are collected below in Lemma 12).
Note that $|g_{r}(x)|=O(1)$ as $r\to\infty$ for any fixed $x\neq 0$ (Lemma 12), while at the origin we have $|g_{r}(0)|=\Omega(r^{(d-1)/2})$, which follows from the series expansion of the Bessel function given in e.g. [Wat80, Section 3.1-3.11]. This causes an issue for $d>1$, as $g_{r}$ is too large at the origin as $r\to\infty$ compared to its tails, which makes it difficult to use it as the difference between two probability densities. Hence, we choose our mollifier $h$ to be supported on an annulus instead of on a ball. In addition, it will also be convenient for it to have a super-polynomially decaying Fourier transform, i.e.
$|\widehat{h}(w)|\leq H(\|w\|)\triangleq C\exp\left(-\frac{c\|w\|}{\log(\|w\|+2)^{2}}\right)\qquad\forall w\in\mathbb{R}^{d}\,.$
The existence of the desired function $h$ is proven Lemma 15.
Note that all of the Fourier energy of $g_{r}$ lies at frequencies $\|\omega\|=r$ by construction. However, after multiplying by $h$ the energy spills over to adjacent frequencies as well and we need to estimate the amount of the spill. Due to the fast decay of $\widehat{h}$ we will show (roughly) the following estimates on the behavior of $\widehat{f}_{r}$:
$|\widehat{f}_{r}(\omega)|$ $\lesssim\tilde{O}(r^{(1-d)/2})1\{\|\omega-r\|\leq\log^{2}(r)\}+r^{(d-1)/2}H(\max(\|\omega\|-r,\log^{2}r)),\quad\text{and}$
$|\widehat{f}_{r}(\omega)|$ $\lesssim r^{(d-1)/2}\|\omega\|$
as $r\to\infty$. Note that the first bound above is super-polynomially decaying in both $\|\omega\|$ and $r$, which allows us to show that
$\|f_{r}\|_{t,2}\leq\tilde{O}(r^{t})\,$
for $t>-\frac{d+2}{2}$, recalling the notation $\|f\|_{t,2}=\|\|\cdot\|^{t}\widehat{f}\|_{2}$. A direct calculation will also show
$\|f_{r}\|_{1}\asymp\|f_{r}\|_{\infty}\asymp\|f_{r}\|_{2}\asymp 1\,.$
For a desired total-variation separation $\epsilon$, we will set $\mu-\nu=\epsilon f_{r}$ and choose $r=\epsilon^{-1/\beta}$ to ensure that $\epsilon\|f_{r}\|_{\beta,2}=\tilde{O}(1)$. For the energy distance between $\mu$ and $\nu$ these choices yield
$\mathcal{E}_{\gamma}(\mu,\nu)\asymp\|\epsilon f_{\epsilon^{-1/\beta}}\|_{-{d+\gamma\over 2},2}=\tilde{O}(\epsilon^{1+\frac{d+\gamma}{2\beta}})=\tilde{O}(\mathsf{TV}^{\frac{d+2\beta+\gamma}{2\beta}})\,,$
as required.
We now proceed to rigorous details.
D.2.2 The construction
First, we must construct the mollifier $h$ with the properties outlined in Section D.2.1. Recall that a function $f$ is radial (a.k.a. spherically symmetric) if its value at $x\in\mathbb{R}^{d}$ depends only on $\|x\|$. In other words, $f(x)=f(y)$ holds for all $x,y\in\mathbb{R}^{d}$ with $\|x\|=\|y\|$.
Lemma 15.
There exists a compactly supported radial Schwartz function $h$, and a positive sequence $\{r_{n}\}_{n=1}^{\infty}$ satisfying $r_{n}=\Theta(n)$, such that
$\operatorname{supp}(h)$ $\subset\mathbb{B}(0,1),$ (D.2)
$\operatorname{supp}(h)$ $\subset\mathbb{R}^{d}\setminus\mathbb{B}(0,r_{0}),$ (D.3)
$|\widehat{h}(w)|$ $\leq C\exp\left(-\frac{c\|w\|}{\log(\|w\|+2)^{2}}\right)\quad\text{for all }w\in\mathbb{R}^{d},\text{ and}$ (D.4)
$\widehat{h}(r_{n}u)$ $=0\quad\text{for all }u\in\mathbb{S}^{d-1},$ (D.5)
for universal constants $C,c,r_{0}>0$.
Proof.
First, let $h_{0}$ be as constructed in Lemma 13, which already satisfies D.2 and D.4. To address the other two requirements, we modify $h_{0}$ by convolving it with two additional terms:
$h(x):=(A_{0}(\cdot)*h_{0}(8\cdot)*\rho_{0}(\cdot))(x),$
where $A_{0}$ and $\rho_{0}$ aim to address D.3 and D.5, respectively, and are defined as
$A_{0}(x)=\exp\left(-\frac{1}{{1/64-(\|x\|-1/2)^{2}}}\right)\mathbbm{1}\big{\{}\|x\|\in(3/8,5/8)\big{\}},\quad\rho_{0}(x)=\mathbbm{1}\{\|x\|<1/8\}.$
Before proceeding, note that clearly $h$ is a radial Schwartz function. Let us now verify that $h$ indeed satisfies the four requirements. Note that $A_{0}$ is an “annulus” supported on $\mathbb{B}(0,5/8)\backslash\mathbb{B}(0,3/8)$, and both $h_{0}(8\cdot)$ and $\rho_{0}$ are supported on $\mathbb{B}(0,1/8)$. Therefore, $\operatorname{supp}(h)\subset\mathbb{B}(0,7/8)\backslash\mathbb{B}(0,1/8)$, which implies D.2 and D.3. We now turn to the other two conditions in Fourier space. Note that
$\widehat{h}(w)=(1/8)^{d}\cdot\widehat{A}_{0}(w)\cdot\widehat{h}_{0}(w/8)\cdot\widehat{\rho}_{0}(w).$
From Item 3 of Lemma 12 we know that
$\mathcal{F}[\mathbbm{1}\{\|\cdot\|<1\}](w)=\left(\frac{2\pi}{\|w\|}\right)^{\frac{d}{2}}J_{\frac{d}{2}}(\|w\|).$
Hence, by Item 2 of Lemma 12, the function $\widehat{\rho}_{0}(w)=(1/8)^{d}\mathcal{F}[\mathbbm{1}\{\|\cdot\|<1\}](w/8)$ has infinitely many zeros near the values of $\|w\|=8(2n\pi+\frac{(d+1)\pi}{4})$ for sufficiently large $n\in\mathbb{Z}^{+}$, which implies D.5.
Finally, for D.4, note that since both $A_{0}$ and $\rho_{0}$ are Schwartz functions, so are their Fourier transforms $\widehat{A}_{0}$ and $\widehat{\rho}_{0}$ so that
$|\widehat{h}(w)|\leq(1/8)^{d}\|\widehat{A}_{0}\|_{\infty}\|\widehat{\rho}_{0}\|_{\infty}|\widehat{h}_{0}(w/8)|\lesssim|\widehat{h}_{0}(w/8)|,$
concluding the proof. ∎
Let $h$ be as constructed in Lemma 15, and define
$f_{r}=g_{r}h$ (D.6)
for $r>0$ and $\widehat{g}(\omega)\triangleq r^{(1-d)/2}\delta(\|\omega\|-r)$. Recall from the overview of our construction that we gave in Section D.2.1 that $f_{r}$ is our proposed density difference which we claim (approximately) saturates Hölder’s inequality in (3.3). The next Lemma records the properties of $f_{r}$ which will enable us to complete our proof.
Lemma 16.
Let $f_{r}$ be as in (D.6) and let $\{r_{n}\}_{n=1}^{\infty}$ be the sequence constructed in Lemma 15. The following hold.
  1. (i)
    For all $n\in\mathbb{N}$ we have
[…]
  2. (ii)
    We have
    $\|f_{r_{n}}\|_{\infty}\asymp\|f_{r_{n}}\|_{2}\asymp\|f_{r_{n}}\|_{1}\asymp 1,$
    hiding constants independent of $n$.
  3. (iii)
    For any $t>-\frac{d+2}{2}$ we have
    $\|f_{r_{n}}\|_{t,2}=O(r_{n}^{t}\log^{d}(r_{n}))$
    as $n\to\infty$, hiding constants independent of $n$.
  4. (iv)
    Recall the definition of $\psi_{\gamma}$ from (2.6). For any $\gamma\in(0,2)$ we have
    $\sup_{v\in\mathbb{S}^{d-1},b\in\mathbb{R}}\left|\int_{\mathbb{R}^{d}}\psi_{\gamma}(\langle x,v\rangle-b)f_{r_{n}}(x)\mathrm{d}x\right|=O(r_{n}^{-(d+\gamma)/2}\log(r_{n})^{d})$
    hiding constants independent of $n$.
Proof.
Let us drop the dependence of $r_{n}$ to simplify notation.
Showing (i). Note that $\int_{\mathbb{R}^{d}}f_{r}(x)\mathrm{d}x=\widehat{f_{r}}(0)$. Then, $\widehat{f_{r}}(0)=0$ follows from the construction of $h$ and $g_{r}$. Indeed, $\widehat{g}_{r}$ is supported on $r\mathbb{S}^{d-1}$ while $\widehat{h}|_{r\mathbb{S}^{d-1}}\equiv 0$. The fact that $\operatorname{supp}(f_{r})\subset\mathbb{B}(0,1)$ follows from $\operatorname{supp}(h)\subset\mathbb{B}(0,1)$.
Showing (ii). Since $f_{r}$ has compact support, we immediately have
$\|f_{r}\|_{1}\lesssim\|f_{r}\|_{2}\lesssim\|f_{r}\|_{\infty}.$
As $h$ is continuous and supported on the annulus $\{x:r_{0}\leq\|x\|\leq 1\}$ by construction, it suffices to bound $g_{r}$ on said annulus. Now, for any $x$ with $r_{0}\leq\|x\|\leq 1$, we have by Lemma 12 that
$g_{r}(x)\lesssim\sqrt{r}\|x\|^{1-d/2}\frac{1}{\sqrt{r\|x\|}}\lesssim 1,$
which shows that $\|f_{r}\|_{\infty}\lesssim 1$.
We now turn to lower bounding $\|f_{r}\|_{1}$. Recall that $h$ is uniformly continuous and nontrivial, hence $\int|h(u^{*}v)|\mathrm{d}\sigma(v)\neq 0$ for some radius $u^{*}$, and thus for all $u\in(u_{0},u_{1})\subseteq(0,1)$ for some constants $u_{0},u_{1}$. Using that $g_{r}$ is spherically symmetric, we compute
$\|f\|_{1}$ $=\int_{\mathbb{R}^{d}}|g_{r}(x)||h(x)|\mathrm{d}x$
$=\int_{0}^{\infty}u^{d-1}g_{r}(u,0,\dots,0)\int h(uv)\mathrm{d}\sigma(v)\mathrm{d}u$
$\gtrsim\sqrt{r}\int_{u_{0}}^{u_{1}}\left|J_{d/2-1}(ru)\right|\mathrm{d}u\gtrsim 1,$
where the last line follows by (A.1) once again.
Showing (iii).  Let $0<s<r$, whose precise value will be set later. For convenience, set $B_{s}=\{x\in\mathbb{R}^{d}:\|x\|\leq s\}$ and $B_{s}^{c}=\mathbb{R}^{d}\setminus B_{s}$. Recall that by definition
$\widehat{f}_{r}(\omega)$ $=r^{(1-d)/2}\int_{\mathbb{R}^{d}}\widehat{h}(\omega+x)\delta(\|x\|-r)\mathrm{d}x$
$=\underbrace{r^{(1-d)/2}\int_{\mathbb{R}^{d}}(\widehat{h}\mathbbm{1}_{B_{s}})(\omega+x)\delta(\|x\|-r)\mathrm{d}x}_{I}+\underbrace{r^{(1-d)/2}\int_{\mathbb{R}^{d}}(\widehat{h}\mathbbm{1}_{B_{s}^{c}})(\omega+x)\delta(\|x\|-r)\mathrm{d}x}_{II}.$ (D.7)
Let $C,c$ be as in Lemma 15, and $H(x)=C\exp(-c\|x\|/\log^{2}(\|x\|+2))$. Note that $\|\widehat{h}\|_{\infty}\leq C$. Therefore, the first term in the decomposition (D.7) can be bounded by
$|I|$ $\leq Cr^{(1-d)/2}\int_{\mathbb{R}^{d}}\mathbbm{1}\{\|\omega+x\|\leq s\}\delta(\|x\|-r)\mathrm{d}x$
$=Cr^{(1-d)/2}\mathbbm{1}\{\|\omega\|\in[r-s,r+s]\}\int_{\mathbb{R}^{d}}\mathbbm{1}\{\|\omega+x\|\leq s\}\delta(\|x\|-r)\mathrm{d}x$
$\lesssim r^{(1-d)/2}s^{d-1}\mathbbm{1}\{\|\omega\|\in[r-s,r+s]\}.$
The second line uses that if $\|\omega\|\not\in[r-s,r+s]$ then the integral becomes zero. The third line uses the fact that the surface area of the intersection of $B_{s}$ with any sphere of any radius (and the one centered at $\omega$ with radius $r$ in particular) is at most $O(s^{d-1})$.
Moving on to the second term, we have
$|II|$ $=r^{(d-1)/2}\int|(\widehat{h}\mathbbm{1}_{B_{s}^{c}})(\omega+ru)|\mathrm{d}\sigma(u)\lesssim r^{(d-1)/2}H(\max\{\|\omega\|-r,s\})$
using that $H:[0,\infty)\to(0,C]$ is decreasing and that $|\widehat{h}(y)\mathbbm{1}_{B_{s}^{c}}(y)\|\leq H(\max\{y,s\})$ for all $y\in\mathbb{R}^{d}$. Summarizing, we have the pointwise estimate
$|\widehat{f}_{r}(\omega)|\lesssim r^{(1-d)/2}s^{d-1}\mathbbm{1}\{\|\omega\|\in[r-s,r+s]\}+r^{(d-1)/2}H(\max\{\|\omega\|-r,s\})$ (D.8)
for all $\omega\in\mathbb{R}^{d}$ and $0<s<r$.
We now show that $f_{r}$ is Lipschitz continuous. Recall from the construction of $h$ (Lemma 15) that $h|_{r_{n}\mathbb{S}^{d-1}}\equiv 0$. Then, we observe that for any $\omega\in\mathbb{R}^{d}$
$|\widehat{f}_{r}(\omega)|$ $=r^{(d-1)/2}\left|\int\widehat{h}(\omega+ru)\mathrm{d}\sigma(u)\right|$
$=r^{(d-1)/2}\left|\int\{\widehat{h}(\omega+ru)-\widehat{h}(ru)\}\mathrm{d}\sigma(u)\right|$
$=r^{(d-1)/2}\|\widehat{h}\|_{\operatorname{Lip}}\frac{2\pi^{d/2}\|\omega\|}{\Gamma(\frac{d}{2})}$
$\lesssim r^{(d-1)/2}\|\omega\|,$ (D.9)
where we use that $\widehat{h}$ is Schwartz by construction, and thus has finite Lipschitz constant $\|\widehat{h}\|_{\operatorname{Lip}}$.
With (D.8) and (D.9) in hand we can proceed to bounding the norm of $f_{r}$. Let $s=D\log(r)^{2}$ for a large constant $D$ independent of $r$, and assume that $r$ is large enough so that $s<r/2$. Also set $\theta>0$, whose precise value is specified later. We have
$\|f_{r}\|_{t,2}^{2}$ $\,\,\,\,=\int_{\mathbb{R}^{d}}\|\omega\|^{2t}|\widehat{f}_{r}(\omega)|^{2}\mathrm{d}\omega$
[…]
$\qquad\qquad+r^{1-d}\log(r)^{2(d-1)}\int_{\|\omega\|>r^{-\theta}}\|\omega\|^{2t}\mathbbm{1}\{\|\omega\|\in[r-s,r+s]\}\mathrm{d}\omega$
$\qquad\qquad+r^{d-1}\int_{\|\omega\|>r^{-\theta}}\|\omega\|^{2t}H^{2}(\max\{\|\omega\|-r,s\})\mathrm{d}\omega$
$\,\,\,\,\lesssim r^{d-1-\theta(2t+d)}+r^{2t}\log(r)^{2d-1}+r^{d-1}\int_{r^{-\theta}}^{\infty}u^{2t+d-1}H^{2}(\max\{u-r,s\})\mathrm{d}u.$
Note that in the derivation above we changed to polar coordinates freely, and that in the second inequality we used the assumption $t>-d/2-1$. Setting $\theta$ to any positive value greater than $(d-1-2t)/(2t+d)$ ensures that the first term in the final line is $O(r^{2t})$. As for the integral term, we can bound it by
[…]
By taking $D$ large enough (independently of $r$) in the definition of $s=D\log^{2}(r)$ we can make also the first term $\operatorname{poly}(r)\times H^{2}(s)$ less than $O(r^{2t})$, which concludes the proof of $(iii)$.
Showing (iv). The bounds that we develop below are analogous to those given in the proof of $(iii)$. Fix $b\in\mathbb{R}$ and $v\in\mathbb{S}^{d-1}$ and define
$\dagger:=\int_{\mathbb{R}^{d}}\psi_{\gamma}(\langle v,x\rangle-b)f_{r}(x)\mathrm{d}{x}.$
Suppose first that $\gamma\neq 1$. Then, using Lemmas 7 and 6, we know by dominated convergence that
$\dagger$ $=\int_{\mathbb{R}^{d}}\lim_{\epsilon\to 0}\int_{\epsilon}^{1/\epsilon}C_{\psi_{\gamma}}\frac{\cos(t(\langle v,x\rangle-b))-\mathbbm{1}\{\gamma>1\}}{t^{(1+\gamma)/2}}f_{r}(x)\mathrm{d}t\mathrm{d}{x}$
$=C_{\psi_{\gamma}}\lim_{\epsilon\to 0}\int_{\epsilon}^{1/\epsilon}\Re\left\{\int_{\mathbb{R}^{d}}\frac{e^{it(\langle v,x\rangle-b)}f_{r}(x)}{t^{(1+\gamma)/2}}\mathrm{d}x\right\}\mathrm{d}t$
$=C_{\psi_{\gamma}}\lim_{\epsilon\to 0}\int_{\epsilon}^{1/\epsilon}\frac{\cos(tb)\widehat{f}_{r}(tv)}{t^{(1+\gamma)/2}}\mathrm{d}t.$
Similarly, for $\gamma=1$ we can compute
$\dagger$ $=C_{\psi_{\gamma}}\lim_{\epsilon\to 0}\int_{\epsilon}^{1/\epsilon}\frac{\sin(-tb)\widehat{f}_{r}(tv)}{t}\mathrm{d}t.$
In either case, we have $|\dagger|\lesssim\int_{0}^{\infty}|\widehat{f}_{r}(tv)|/t^{(1+\gamma)/2}\mathrm{d}t$.
Let $s=D\log^{2}(r)$ for large $D$ independent of $r$ as in the proof of $(iii)$, and let $\theta>0$ whose precise value is specified later. Assuming that $r$ is large enough so that $s<r/2$, for any $\gamma\in(0,2)$ we have
||$\dagger$|| † | $\,\,\,\,\leq\int_{0}^{r^{-\theta}}\frac{|\widehat{f}_{r}(tv)|}{t^{(1+\gamma)/2}}\mathrm{d}t+\int_{r^{-\theta}}^{\infty}\frac{|\widehat{f}_{r}(tv)|}{t^{(1+\gamma)/2}}\mathrm{d}t$
[…]
$\qquad+\int_{r^{-\theta}}^{\infty}\frac{1}{t^{(1+\gamma)/2}}\left(r^{(1-d)/2}s^{d-1}\mathbbm{1}\{t\in[r-s,r+s]\}+r^{(d-1)/2}H(\max\{t-r,s\})\right)\mathrm{d}t$
$\,\,\,\,\lesssim r^{\frac{d-1}{2}-\theta\frac{3-\gamma}{2}}+r^{-(d+\gamma)/2}\log^{d}(r)+H(s)r^{(d-1)/2}\int_{r^{-\theta}}^{2r}\frac{\mathrm{d}t}{t^{(1+\gamma)/2}}+\int_{2r}^{\infty}\frac{H(t/2)}{t^{(1+\gamma)/2}}\mathrm{d}t$
[…]
Set $\theta$ to any value greater than $(2d+\gamma-1)/(3-\gamma)$, which ensures that the first term in (D.10) is $O(r^{-(d+\gamma)/2})$. By taking $D$ large enough in the definition of $s=D\log^{2}(r)$, we can make $H(s)$ smaller than any polynomial in $r$, which ensures that the third term in (D.10) is also $O(r^{-(d+\gamma)/2})$. We thus obtain the final bound $|\dagger|\lesssim r^{-(d+\gamma)/2}\log^{d}(r)$, concluding our proof. ∎
Proof of Theorem 7 and Proposition 8 for $d>1$.
Using the functions $\{f_{r_{n}}\}_{n=1}^{\infty}$ we constructed in Lemma 16, we are ready to prove Propositions 7 and 8 for $d>1$.
Let $p_{0}$ be a compactly supported probability density with $\inf_{\|x\|\leq 1}p_{0}(x)>0$. Fix the smoothness $\beta>0$. Given any desired total variation separation $\epsilon\in(0,1)$, we can find $n_{0}\in\mathbb{N}$ such that $\epsilon^{-1/\beta}\asymp r_{n_{0}}$, where we hide an $\epsilon$-independent multiplicative constant. Define
$p_{\epsilon}=p_{0}+\epsilon f_{r_{n_{0}}}/2\qquad\text{and}\qquad q_{\epsilon}=p_{0}-\epsilon f_{r_{n_{0}}}/2.$
Clearly $p_{\epsilon}$ and $q_{\epsilon}$ are compactly supported probability densities for all small enough $\epsilon$. Moreover, by Lemma 16 they satisfy
$\|p_{\epsilon}-q_{\epsilon}\|_{1}\asymp\epsilon$ $\qquad\|p_{\epsilon}\|_{\beta,2}\asymp\|q_{\epsilon}\|_{\beta,2}\asymp 1\qquad\text{and}$ and
$\mathcal{E}_{\gamma}(p_{\epsilon},q_{\epsilon})\lesssim\epsilon^{\frac{2\beta+d+\gamma}{2\beta}}\log(1/\epsilon)^{d}$ $\qquad\text{and}\qquad\overline{d_{H}}(p_{\epsilon},q_{\epsilon})\lesssim\epsilon^{\frac{2\beta+d+1}{2\beta}}\log(1/\epsilon)^{d}$
for all fixed $\gamma\in(0,2)$. This concludes our proof. ∎
Appendix E Proof of Proposition 10
Proof.
Note that $\overline{d_{H}}=T_{d,0}$. Let $p_{\epsilon},q_{\epsilon}$ be the compactly supported densities constructed in the proof of Proposition 7 in the general dimensional case. Then by construction
$\varepsilon\asymp\mathsf{TV}(p_{\epsilon},q_{\epsilon})\asymp\|p_{\epsilon}-q_{\epsilon}\|_{2}\quad\text{and}\quad\|p_{\epsilon}\|_{\beta,2}+\|q_{\epsilon}\|_{\beta,2}\lesssim 1\quad\text{and}\quad\overline{d_{H}}(p_{\epsilon},q_{\epsilon})\lesssim\epsilon^{\frac{2\beta+d+1}{\beta}}\log(1/\epsilon)^{d}.$
Write $p_{\epsilon,n}$ and $q_{\epsilon,n}$ for the empirical measures of $p_{\epsilon}$ and $q_{\epsilon}$ respectively, based on $n$ i.i.d. observations each. By the triangle inequality we have
$\mathbb{E}\overline{d_{H}}(p_{\epsilon,n},q_{\epsilon,n})$ $\leq\mathbb{E}\overline{d_{H}}(p_{\epsilon,n},p_{\epsilon})+\overline{d_{H}}(p_{\epsilon},q_{\epsilon})+\mathbb{E}\overline{d_{H}}(q_{\epsilon},q_{\epsilon,n})$
[…]
This completes the proof. ∎